### v1

# Benchmarking Computer Science Survey Generation

Abstract
Scientific survey articles play a vital role in summarizing research progress, yet their manual creation is becoming increasingly infeasible due to the rapid growth of academic literature. While large language models (LLMs) offer promising capabilities for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols. To address this gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for evaluating scientific survey generation in the computer science domain. SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers that serves as the retrieval pool. In addition, we propose an automated evaluation framework that measures generated surveys across four dimensions: information coverage, referencing accuracy, structural organization, and content quality. Our evaluation of diverse LLM-based approaches shows that survey generation remains highly challenging, even for advanced self-reflection frameworks. These findings highlight the complexity of the task and the necessity for continued research [^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
1 Introduction
Large language models (LLMs) have shown impressive capabilities in scientific text generation wang2024autosurvey ; gao2024evaluating ; liang2025surveyx , supporting various stages of academic writing from outlining to full-content generation. Among these applications, automated scientific survey generation stands out as a particularly valuable task. Scientific survey articles (also known as survey papers or literature reviews) synthesize knowledge from a broad collection of academic papers into a structured overview of a research area keele2007guidelines ; pautasso2013ten ; boote2005scholars . Traditionally, writing such surveys is highly labor-intensive: an expert in this field must gather dozens or even hundreds of relevant papers, understand their contributions, identify important themes and trends, and finally craft a well-structured survey paper. This labor-intensive process has become increasingly challenging due to the explosive growth of academic literature in recent years. For example, arXiv now receives over a thousand new computer science papers each day, and the annual volume of submissions has more than doubled from 2019 to 2024 liang2025surveyx . Keeping pace with this scale of information is infeasible for individual researchers, prompting growing interest in automatic survey writing.
Despite the success of LLMs in various generation tasks, generating structured and comprehensive scientific surveys has received relatively little attention. A key reason is the lack of standardized benchmarks and evaluation protocols tailored to the unique demands of this task. Unlike general-purpose generation tasks, scientific survey writing remains difficult to evaluate in a quantitative and standardized way, due to the lack of benchmarks that capture its unique requirements: synthesizing multiple sources, ensuring citation accuracy, and maintaining structural coherence. As a result, current systems are often assessed through manual expert reviews nlpcc2024 , which are costly, time-consuming, and hinder reproducibility and systematic progress.
To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark designed to facilitate reproducible research in scientific survey generation. We formalize the task as a two-stage process: given a topic description in the computer science domain (e.g., Machine Learning for Information Retrieval), the system should first collect a set of relevant papers from a large academic corpus and then synthesize a well-structured survey from them. To support this formulation, SurGE provides the necessary resources and a standardized evaluation framework for both the retrieval and generation stages. Each test instance in SurGE comprises a research topic and its corresponding ground-truth survey, which serves as a high-quality reference for evaluation. These ground-truth surveys are carefully selected from peer-reviewed, high-impact survey papers, based on citation statistics and topical relevance. SurGE dataset consists of two main components: (1) a set of test instances, where each instance includes a topic description, its ground-truth survey, and the full list of cited references in that survey; (2) a large-scale academic corpus containing over one million computer science papers, which serves as the retrieval pool for relevant document selection and coverage analysis.
Beyond dataset construction, we also propose an automated evaluation framework that enables reproducible and fine-grained assessment of generated surveys. This framework scores outputs across four key dimensions: information collection, referencing accuracy, structural organization, and content quality. Grounded in established academic writing guidelines webster2002analyzing ; boote2005scholars ; keele2007guidelines ; pautasso2013ten , this framework enables automatic, scalable evaluation without the need for human annotation, making it well-suited for benchmarking and driving systematic progress in this emerging task.
To validate the utility of SurGE, we evaluate a range of LLM-based baseline systems on the benchmark. These include standard retrieval-augmented generation approaches that first retrieve relevant documents and then generate a survey, as well as more advanced pipelines that incorporate explicit planning (e.g., outline generation) and iterative refinement of the draft. The results reveal that even the state-of-the-art systems struggle with the survey generation task, highlighting its difficulty. For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant references. These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research. We hope that SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
In summary, our contributions are threefold:
  1. 1.
    We introduce SurGE, the first comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
  2. 2.
    We propose a fully automated evaluation framework that assesses survey quality across four crucial dimensions: information coverage, citation accuracy, structure, and content.
  3. 3.
    We provide extensive baseline results and analyses, offering reference points and highlighting key challenges in this emerging task.
2 Task Definition
We formalize scientific survey generation as a two-stage task. Given a topic description $t$ and a large academic corpus $D=\{d_{1},d_{2},\ldots,d_{n}\}$, the goal is to automatically generate a survey article $S$ that provides a structured and comprehensive overview of the topic. The process consists of:
  • Document Collection: A retrieval module or complex retrieval system identifies a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
  • Survey Generation: A generative model composes a well-structured survey $S$ based on the topic $t$ and the retrieved document set $\mathcal{R}_{t}$, including proper citations and a reference list.
The output survey $S$ is expected to (i) follow a logical hierarchical organization, (ii) cover key techniques and trends related to the topic, and (iii) provide accurate citations in the appropriate places using a consistent and correct format.
3 Dataset Construction
3.1 Ground-Truth Survey Collection and Annotation
To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature. Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews. To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact bornmann2008citation ; and (iii) the publish date must be between 2020 and 2024.
Following the initial filtering process, we further refined the SurGE dataset through expert annotation to ensure the inclusion of only high-quality survey documents. This stage aimed to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers. To this end, we recruited a team of four computer science Ph.D. students as annotators. Note that annotating a paper did not require a close reading of the entire document. Annotators could typically complete the task within 8–9 minutes per paper on average. Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) reference quality, which examines the relevance, diversity, and traceability of cited works.
Each annotator labeled the document as either “usable” or “not usable.” A survey paper was included in the final dataset only if both annotators independently marked it as usable. In cases of disagreement, the paper was discarded to maintain a conservative quality threshold. To ensure fair and motivated participation, annotators were compensated based on working hours at an average rate of 60 CNY per hour, exceeding the local minimum wage in Beijing. Inter-annotator agreement was quantified using Cohen’s Kappa, applied to 250 annotated instances. The resulting score of 0.792 indicates substantial agreement and reinforces the reliability of the quality control process. After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers. The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix C.
3.2 Reference Extraction and Processing
[…]
The core reason for extracting the full reference list from each ground-truth survey is to create a gold standard for evaluating the information collection stage. We operate on the premise that the citations in an expert-written survey represent a curated collection of the field’s most foundational and relevant literature. By using this set as the ground truth, we can quantitatively measure the coverage and recall of a system-generated survey’s references, providing a clear metric for its performance.
3.3 Academic Corpus Construction
A crucial component of the SurGE benchmark is a large-scale academic corpus designed to serve as the retrieval pool for the document collection stage. Our corpus is built entirely from scholarly metadata obtained from the arXiv repository. To ensure adherence to ethical and legal standards, we exclusively collected metadata and did not include full-text PDFs, a practice permitted by arXiv’s Terms of Use, which designates metadata as public domain under the CC0 license arxiv_tou .
The corpus was constructed through a two-stage process. The initial stage involved seeding the corpus with the references from our 205 ground-truth surveys. We systematically retrieved the arXiv metadata for every cited paper that was publicly accessible. This process revealed that approximately 30% of the references were unavailable, primarily due to publication in closed-access journals or other restricted venues. In the second stage, we expanded the corpus to enhance its comprehensiveness. We queried the official arXiv search API, using keywords and titles from the ground-truth surveys to identify and collect metadata for other topically related papers. This methodology resulted in a final retrieval corpus of 1,086,992 unique papers. For each paper, the corpus provides rich metadata, including its title, authors, abstract, subject categories, publication date, and a direct link to its arXiv page for transparency and verification. To ensure high data quality, all collected metadata underwent a rigorous pre-processing pipeline involving text normalization, formatting removal, and deduplication.
3.4 Statistics and Analysis
The resulting SurGE benchmark comprises 205 ground-truth survey papers and a retrieval corpus of 1,086,992 documents. Table 1(a) presents the key statistics of our curated dataset. To quantitatively analyze the organizational complexity of the surveys, we model the hierarchical section headings (e.g., Section, Subsection) of each paper as a tree structure. Our analysis reveals that these surveys are structurally deep, with an average tree depth of 3.07 and a mean of 42.7 nodes (i.e., distinct sections) per document. This structural complexity presents a significant challenge for hierarchical text generation. Furthermore, the surveys are densely referenced, citing an average of 65.8 papers, which underscores the demand for high-recall information collection. Table 1(b) details the pre-processed fields for each survey instance, which include not only standard metadata but also the parsed structural tree and the ground-truth list of cited documents, enabling a fine-grained, multi-faceted evaluation of system-generated surveys.
3.5 Ethical Considerations and Licensing
Our corpus is constructed exclusively from arXiv-provided descriptive metadata (titles, authors, abstracts, identifiers, categories, and license URIs) harvested via the official API. We do not host or redistribute arXiv PDFs or source files. This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication arxiv_tou . This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata arxiv_submission .
We have prioritized transparency, reproducibility, and ethical considerations throughout dataset construction. To support open science, we have publicly released the SurGE dataset, accompanying metadata, and all associated processing scripts on our official GitHub repository [^2] https://github.com/oneal2000/SurGE. The dataset and codebase are distributed under the MIT license, granting researchers and developers unrestricted access and modification rights. Regular updates will ensure continued relevance and alignment with evolving research trends and standards.
(a) Basic Statistics of the SurGE Dataset and Corpus
Statistic Number
Total Ground Truth Surveys 205
Average Tree Depth 3.073
Maximum Tree Depth 4
Average Number of Tree Nodes 42.717
Maximum Number of Tree Nodes 212
Average Citations per Paper 65.78
Average Citations per Section 1.577
Corpus Size 1,086,992
Average Abstract Length (words) 156.57
(b) Fields in the Pre-processed Survey Dataset
Field Description
SurveyID A unique identifier for the survey.
Authors List of contributing researchers.
Title The title of the survey paper.
Year The publication year of the survey.
Date The timestamp of publication.
Category Subject classification following the arXiv taxonomy.
Abstract The abstract of the survey paper.
Structure Hierarchical representation of the survey.
All_Cites List of document IDs cited in the survey.
Figure 1: Overview of the SurGE Benchmark. (a) Summary statistics of the curated survey dataset and its associated retrieval corpus. (b) Metadata of the pre-processed survey dataset used in SurGE.
4 Evaluation Framework
Developing a robust evaluation framework for automated scientific survey generation involves identifying critical dimensions that reflect a survey’s overall accuracy, comprehensiveness, and quality. Through discussions with academic experts, peer reviewers, and experienced survey authors, we established four essential evaluation dimensions:Information Collection Quality, Referencing Accuracy, Structural Quality, and Content Quality. Each generated survey is systematically compared with an expert-written Ground Truth (GT) survey, ensuring alignment with high academic standards. In the following subsections, we deeply analyze each evaluation criterion and subsequently define the corresponding quantitative metrics.
4.1 Information Collection Quality
A fundamental requirement for scientific surveys is the comprehensive collection of relevant literature, covering key works that accurately represent the research field. Omissions or insufficient coverage can severely compromise a survey’s credibility, causing readers to question its thoroughness. Therefore, this aspect is quantitatively assessed through the Coverage Score, measuring the extent to which a generated survey includes key literature identified by expert consensus in the GT survey.
Metric Definition.
The coverage score $\mathcal{C}$ is defined as the recall of references included in the generated survey relative to the GT survey:
$\mathcal{C}=\frac{|{R}_{GT}\cap{R}_{G}|}{|{R}_{GT}|},$ (1)
where ${R}_{GT}$ denotes references in the GT survey, and ${R}_{G}$ denotes references in the generated survey. A higher coverage score indicates more comprehensive inclusion of essential literature.
While the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited, peer-reviewed publications. We therefore interpret the Coverage score not as a measure of absolute completeness, but as a pragmatic metric for evaluating a system’s ability to identify the central body of work validated by the research community.
4.2 Referencing Accuracy
Referencing Accuracy evaluates how well the references cited in a generated survey align with its topic and content. This assessment is crucial for ensuring that the survey is well-supported and that citations are contextually appropriate. We propose a hierarchical evaluation framework that assesses referencing quality from three perspectives:
  • Document-level: Determines if a cited paper is thematically relevant to the survey’s overall topic.
  • Section-level: Evaluates if a reference is placed within a semantically appropriate section of the survey.
  • Sentence-level: Assesses if a reference directly supports the specific claim made in the sentence where it is cited.
4.2.1 Automated Relevance Judgment.
To automate this evaluation, we employ a Natural Language Inference (NLI) model (cross-encoder/nli-deberta-v3-base) to automatically assess relevance. For each reference, we form a premise-hypothesis pair and use the NLI model to predict the relationship between them. The premise is constructed from the reference’s metadata, while the hypothesis is contextualized by its placement in the generated survey.
Specifically, a reference is judged based on the NLI model’s output scores for “Entailment,” “Neutral,” and “Contradiction”:
  • It is deemed relevant if “Entailment” has the highest score.
  • It is deemed partially relevant if “Neutral” is highest and “Entailment” is the second-highest.
  • Otherwise, it is deemed irrelevant.
Since the survey generation process is grounded in our academic corpus, any reference cited by the model that is not found within this corpus is considered a hallucination. Such references are therefore automatically marked as irrelevant across all three hierarchical levels.
4.2.2 Hierarchical Evaluation Metrics.
1. Document-level Accuracy ($R_{d}$). This metric assesses if each reference is relevant to the survey’s main topic. For a cited reference $r$, we first apply a heuristic: if $r$ is also cited in the ground-truth survey, it is automatically marked as relevant. For all other references, we use the NLI model with the following pair:
  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract could be cited in the paper: ‘{Survey Title}’.”
2. Section-level Accuracy ($R_{s}$). This metric evaluates the contextual appropriateness of a citation within a specific section. We assess each unique (reference, section) pair. If a citation marker in the text points to a non-existent entry in the bibliography, the pair is marked as irrelevant. Otherwise, we query the NLI model:
  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract is relevant to the section: ‘{Section Title}’.”
3. Sentence-level Accuracy ($R_{t}$). For the most fine-grained evaluation, this metric measures if a reference directly supports the claim in its citing sentence. We assess each unique (reference, sentence) pair using the NLI model:
  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract could be cited in the sentence: ‘{Sentence Text}’.”
4.2.3 Metric Calculation.
Based on the automated judgments, we define indicator functions for each reference $r$ at a given level: $\mathbb{I}_{\text{rel}}(r)=1$ if $r$ is relevant, and 0 otherwise; $\mathbb{I}_{\text{half}}(r)=1$ if $r$ is partially relevant, and 0 otherwise. The final accuracy score for each level $x\in\{d,s,t\}$ is calculated as:
$R_{x}=\frac{\left|\{r\in\mathcal{R}_{x}\mid r\text{ is relevant}\}\right|+0.5\cdot\left|\{r\in\mathcal{R}_{x}\mid r\text{ is partially relevant}\}\right|}{\left|\mathcal{R}_{x}\right|},$ (2)
where $\mathcal{R}_{d}$ is the set of all unique references in the generated survey, while $\mathcal{R}_{s}$ and $\mathcal{R}_{t}$ are the sets of all unique (reference, section) and (reference, sentence) citation instances, respectively.
4.3 Structural Quality
A clear and logical structure is fundamental to a high-quality scientific survey, as it greatly enhances readability, coherence, and the ability to quickly navigate information. In contrast, poor structural organization can cause confusion, redundancy, and difficulty in following the paper’s arguments. Therefore, a robust evaluation of structural quality is essential. We propose to assess structure from two complementary perspectives: coarse-grained, high-level organization and fine-grained, heading-level alignment.
To this end, we introduce two metrics for structural assessment:
  • Structure Quality Score (SQS): Assesses the high-level structural alignment by comparing the list of generated headings with the ground-truth (GT) headings, focusing on their similarity in structure, meaning, and wording.
  • Soft-Heading Recall (SHR): Measures the fine-grained alignment of headings and subheadings by using semantic embeddings to account for partial or paraphrased matches.
Structure Quality Score (SQS).
SQS is designed to evaluate the overall quality of the generated survey’s outline. For this evaluation, we define an outline as a hierarchical list of section and subsection headings extracted from a survey. SQS assesses the high-level structural alignment by comparing the generated outline against the ground-truth (GT) outline, focusing on aspects like topical coverage, logical flow, and hierarchical organization. To compute this score, we employ the "LLM-as-a-Judge" paradigm. [^3] The corresponding prompt templates are available in our official GitHub repository: https://github.com/oneal2000/SurGE. Specifically, we prompt a powerful large language model, GPT-4o, with both the generated and the GT outlines. The model is then instructed to provide a holistic quality score from 1 (poor) to 5 (excellent).
Soft-Heading Recall (SHR).
To measure fine-grained alignment, SHR evaluates how well the generated outline covers the specific headings present in the ground-truth outline. Unlike metrics based on exact lexical matching, SHR leverages semantic similarity to robustly handle variations in wording and paraphrasing.
Formally, SHR is defined as the soft cardinality overlap between the predicted heading set ($H_{P}$) and the ground-truth heading set ($H_{GT}$):
SHR $=\frac{\mathcal{S}(H_{P}\cap H_{GT})}{\mathcal{S}(H_{GT})}$ (3)
where $\mathcal{S}(A)$ denotes the soft cardinality of a heading set $A$. It is designed to down-weight redundant headings within the same set and is calculated as:
$\mathcal{S}(A)$ $=\sum_{i=1}^{K}\frac{1}{\sum_{j=1}^{K}\text{sim}(A_{i},A_{j})}$ (4)
Here, $\text{sim}(A_{i},A_{j})$ is the cosine similarity between the embeddings of headings $A_{i}$ and $A_{j}$:
$\text{sim}(A_{i},A_{j})$ $=\cos(\text{embed}(A_{i}),\text{embed}(A_{j}))$ (5)
Since a standard set intersection would be too strict for comparing paraphrased headings, we calculate the soft cardinality of the intersection $\mathcal{S}(H_{P}\cap H_{GT})$ using the soft set inclusion-exclusion principle:
$\mathcal{S}(H_{P}\cap H_{GT})=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})-\mathcal{S}(H_{P}\cup H_{GT})$ (6)
This formulation ensures that semantically matched headings contribute positively to the recall, even with lexical variations. A higher SHR score indicates better structural alignment with the GT outline at a granular level.
4.4 Content Quality
Content quality evaluation is pivotal for ensuring surveys are not only comprehensive but also clear, coherent, and logically sound. Evaluating this dimension involves assessing the presence of essential information, accuracy, and fluency of expression. While traditional metrics such as ROUGE and BLEU effectively measure textual overlap and fidelity, they inadequately capture nuanced aspects of logical coherence and clarity. To address these limitations, we supplement these traditional metrics with a comprehensive logical coherence evaluation.
Metric Definitions.
We apply ROUGE and BLEU at the section-level to effectively capture fidelity and precision in content representation without needing explicit mathematical formulations here. Logical coherence is evaluated using GPT-4o to assign readability and coherence scores (0–5) for each section.  [^4] Here, we employ the LLM-as-a-Judge paradigm for evaluation. The corresponding prompt templates are available on our official GitHub repository: https://github.com/oneal2000/SurGE.
5 Baseline Implementation
In this section, we detail the implementation of baseline methods. Each baseline follows a two-stage pipeline: (1) retrieving a set of potentially relevant papers for a given topic, and (2) organizing and summarizing the retrieved content to produce a structured survey. To ensure a fair comparison, all baselines rely on a shared dense retriever for the first stage. In the following subsections, we first describe the training procedure of the shared Paper Retriever, followed by the specific methodologies adopted by each baseline in generating the final survey.
5.1 Paper Retriever Training
We employ a dual-encoder architecture for document retrieval, initialized with roberta-base. For a query $q$ and a paper abstract $d$, we first construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$. Formally, let $X(q)=[\mathrm{CLS}]\,q\,[\mathrm{SEP}]$ and $X(d)=[\mathrm{CLS}]\,d\,[\mathrm{SEP}]$. We then feed these tokens into the transformer encoder to obtain a hidden vector at the $\mathrm{CLS}]$ position:
$\mathrm{Emb}(X)=\mathrm{transformer}_{[\mathrm{CLS}]}(X).$ (7)
The similarity score between the query and the document is computed as the dot product of their embeddings:
$S(q,d)=\mathrm{Emb}(X(q))^{\top}\cdot\mathrm{Emb}(X(d)).$ (8)
During training, each query $Q$ is paired with a relevant document $d^{+}$ to form a positive sample, while negative samples $d^{-}\in N$ are drawn from the corpus. The retriever is optimized via the softmax cross-entropy loss:
$\mathcal{L}(Q,d^{+},N)=-\log\frac{\exp\bigl{(}S(Q,d^{+})\bigr{)}}{\exp\bigl{(}S(Q,d^{+})\bigr{)}+\sum_{d^{-}\in N}\exp\bigl{(}S(Q,d^{-})\bigr{)}}.$ (9)
This objective encourages the model to assign higher scores to relevant papers while minimizing scores for irrelevant ones. We use the trained retriever to retrieve the top-ranked papers for each query, thereby providing a focused subset of literature for subsequent summarization.
5.2 Retrieval-Augmented Generation
In the first baseline, we combine retrieval with a direct generation approach. Given a topic $t$, we use the above retriever to collect the top 100 candidate papers. To manage lengthy inputs, these retrieved papers are split into smaller groups, each containing an approximately equal number of references. We then prompt a large language model (LLM) to summarize each group separately, guiding it to preserve references to the original papers. Formally, for each group of papers $G_{k}=\{d_{1},d_{2},\ldots,d_{n_{k}}\}$, the LLM is conditioned on the sequence of paper abstracts and instructed to produce a partial summary $\hat{S}_{k}$. Finally, we merge the partial summaries $\hat{S}_{1},\hat{S}_{2},\ldots$ into a unified survey. This merging step is performed by prompting the LLM once again with all partial summaries and asking for an integrated, logically coherent survey. Although this baseline follows a straightforward two-step approach, it provides a clear assessment of how effective retrieval-based summarization can be when coupled with an LLM’s generative capabilities.
5.3 AutoSurvey
AutoSurvey wang2024autosurvey implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions. We adapt it to use our fine-tuned retriever in place of its original retrieval mechanism and keep the number of retrieved references consistent for fairness. In the adapted workflow, we first issue a query based on the topic $t$ to retrieve an initial collection of papers $P_{\text{init}}$. The LLM is then prompted to create a structured outline, which includes main sections and subsections tailored to the subject matter. Next, each section is expanded by conditioning on the subset of papers most relevant to that specific section, producing a draft that includes references in a bracketed format (e.g., “[id]”). Once each section is drafted, the LLM refines it to address factual inconsistencies, stylistic mismatches, and reference-formatting issues. Finally, all refined sections are merged into a coherent final survey, with transitions and citation references carefully aligned. The workflow iterates over these stages, leading to incremental improvements in both thematic coverage and presentation quality.
5.4 StepSurvey
StepSurvey lai2024instruct is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner. This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 tian2024overview . Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature. Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers. The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections. Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy. By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
5.5 Implementation Details.
For the training of Paper Retriever, we randomly split the dataset into a training set and a test set at a ratio of 4:1. We adopt the AdamW optimizer for model optimization, the learning rate is set to $5\times 10^{-6}$, the weight decay is set to $1\times 10^{-2}$, and the epoch is set to $10$. During the training process, we adopt mixed-precision (fp16) training for enhanced efficiency. At inference time, each query retrieves the top 100 most relevant papers according to the similarity score $S(q,d)$. The retriever is initialized using the pre-trained RoBERTa model liu2019roberta . For the generation configuration of LLMs, all experiments are conducted using the publicly available implementations provided by Hugging Face. We utilize the default hyperparameters and the chat template as outlined in the official Hugging Face repository  [^5] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct. All experiments are conducted on a server with 1TB of RAM and eight NVIDIA A100 GPUs, each with 40GB of memory.
6 Experimental Results
This section demonstrates the utility of our proposed benchmark by presenting a detailed evaluation of three representative survey generation baselines: RAG, AutoSurvey, and StepSurvey. Our analysis is designed to showcase how the benchmark’s multi-faceted metrics can be used to conduct a fine-grained analysis and pinpoint system bottlenecks. We first isolate and evaluate the performance of the shared Paper Retriever (§6.1). This analysis is crucial as it establishes the theoretical upper-bound performance for the reference coverage metric and highlights the challenges posed by the initial retrieval stage. Following this, we assess the complete end-to-end performance of the baselines (§6.2), leveraging our comprehensive evaluation suite to compare their architectural strengths and weaknesses and to shed light on the broader challenges facing automatic survey generation.
Table 1: Comparison of retrieval models on recalling ground-truth cited papers. The metric is Recall@k, where $k$ is the number of top documents retrieved. Best results are in bold.
Model Recall@20 Recall@30 Recall@100 Recall@200 Recall@500 Recall@1000
BM25 0.0548 0.0652 0.1193 0.1596 0.2213 0.2715
Paper Retriever 0.1706 0.2145 0.3665 0.4681 0.6011 0.6805
6.1 Analysis of Retrieval Performance
A crucial question in our two-stage pipeline is whether performance limitations stem from the retriever’s inability to find relevant papers or the generator’s inability to use them. To disentangle these factors and quantify the retrieval bottleneck, we first evaluate the performance of our fine-tuned dense retriever in isolation. This analysis establishes the theoretical upper bound for the reference coverage that our end-to-end systems can achieve.
We compare our dense retriever against a lexical baseline BM25 robertson2009probabilistic , using Recall@k as the evaluation metric. This metric measures the percentage of ground-truth papers from the reference survey that are present in the top-k retrieved documents. As shown in Table 1, our fine-tuned Paper Retriever substantially outperforms the BM25 baseline across all values of k. The performance gap underscores the inadequacy of lexical search, which struggles to find semantically relevant papers that may not share overlapping keywords. Our dense retriever, by capturing deeper semantic relationships, is far more effective. However, the results also reveal a critical bottleneck in the survey generation pipeline. Even with a generous budget of k=1000, Paper Retriever recalls only 68.05% of the ground-truth papers. This finding is crucial as it establishes a hard upper-bound on the performance of the downstream generator. This finding suggests that achieving a breakthrough in survey quality may necessitate a shift towards more sophisticated retrieval paradigms, such as employing search agents powered by large language models zhang2024large .
Table 2: The experimental results combining Coverage, Referencing Accuracy (including Doc-Rel, Sec-Rel, and Sent-Rel), Structural Quality (including SQS and SHR), and Content Quality (as measured by ROUGE scores, BLEU, and Logic Score). Doc-Rel represents document-level relevance, Sec-Rel represents section-level relevance, Sent-Rel represents sentence-level relevance, SQS represents the structure quality score, and SHR represents soft-heading recall. R-L represents ROUGE-L. The best results are in bold, and the second-best results are underlined.
Coverage Referencing Accuracy Structural Quality Content Quality
Baseline Recall Doc-Rel Sec-Rel Sent-Rel SQS SHR R-L BLEU Logic Score
RAG 0.0214 0.2857 0.2502 0.2500 0.6829 0.7900 0.1519 10.38 4.6723
AutoSurvey 0.0351 0.3617 0.4935 0.4870 1.3902 0.9697 0.1578 10.44 4.7390
StepSruvey 0.0630 0.4576 0.4571 0.4636 1.1951 0.9763 0.1590 12.02 4.8451
6.2 Overall Performance of Survey Generation Baselines
The preceding analysis revealed that the retriever provides the generation models with over a third (36.65%) of the ground-truth references, setting a clear performance ceiling. We now demonstrate the full diagnostic power of our benchmark by applying its comprehensive suite of metrics to the end-to-end outputs of the three baselines: RAG, AutoSurvey, and StepSurvey. As detailed in Table 2, this multi-faceted evaluation allows us to pinpoint specific challenges in the generation stage. The key insights revealed by our benchmark are as follows:
(1) Bottleneck of Existing Baselines. A striking observation across all methods is the significant gap between retrieval performance and final survey coverage. While our paper retriever makes 36.65% of the ground-truth references available (Recall@100), the best-performing baseline, StepSurvey, only achieves a final Coverage of 6.30%. The simpler RAG baseline is even lower at 2.14%. This disparity underscores that the primary limitation is the generation stage’s inability to effectively identify and incorporate relevant information from the provided set of 100 papers. Nonetheless, the results also show that advanced, structured approaches like AutoSurvey (3.51%) and StepSurvey (6.30%) are substantially better at utilizing the retrieved context than a standard RAG pipeline.
(2) Strengths of AutoSurvey. AutoSurvey exhibits excellent Section-Level Relevance (0.4935) and Sentence-Level Relevance (0.4870), surpassing both RAG and StepSurvey. This improvement appears to stem from AutoSurvey’s iterative planning mechanism, where the system generates a high-level outline, retrieves relevant references for each section, and incrementally refines each subsection. By focusing on context-specific references and refining drafts section by section, AutoSurvey places citations more accurately at both the section and sentence levels. Moreover, AutoSurvey achieves the highest Structure Quality Score (SQS) of 1.3902, indicating that its top-down planning leads to well-organized hierarchical structures.
(3) Strengths of StepSurvey. On the other hand, StepSurvey attains the best performance in several crucial dimensions: Coverage (6.30%) and Document-Level Relevance (0.4576). Its multi-phase workflow first generates major headings, then subtopics, and finally assigns references to each subsection. This step-by-step refinement appears particularly effective at broadening the coverage of included works and aligning the cited literature with the overarching survey theme. StepSurvey also achieves the highest Logic Score (4.8451) and excels in ROUGE and BLEU, reflecting strong content quality and coherent presentation. An interesting observation is that while StepSurvey’s Section-Level Relevance is slightly lower than AutoSurvey’s, the difference remains small (0.4571 vs. 0.4935), indicating that its primary strength lies in more holistic, document-wide alignment and consistent logical flow.
(4) Impact of Planning on Multi-Level Relevance. The multi-stage generation process adopted by AutoSurvey and StepSurvey benefits Document-Level and Section-Level Relevance in different ways. AutoSurvey’s approach—outlining, drafting, and iteratively refining sections—yields highly relevant citations at the subsection and sentence granularity. In contrast, StepSurvey’s method of organizing subtopics around central themes ensures strong document-wide alignment. This discrepancy suggests that carefully balancing local vs. global planning remains an open challenge: future work might explore hybrid strategies that combine AutoSurvey’s fine-grained focus with StepSurvey’s broad thematic coverage.
In conclusion, our results demonstrate that advanced, multi-stage planning approaches (AutoSurvey and StepSurvey) significantly outperform the standard RAG pipeline in nearly all metrics. While AutoSurvey excels in section- and sentence-level citation placement and high-level structural quality, StepSurvey provides superior document-level coverage, global relevance, and strong logical coherence. Nonetheless, all methods still struggle to capture the vast reference space fully. These findings underscore the importance of iterative refinement and structured planning in automated survey generation and suggest that there remains substantial room for further improvement, particularly in boosting coverage and refining the interplay between local and global survey organizations.
7 Related Work
7.1 Retrieval-Augmented Generation
Large Language Models (LLMs) are inherently limited by their static, pre-trained parametric knowledge. To address these limitations, Retrieval-Augmented Generation (RAG) has emerged as a key paradigm gao2023retrieval ; lewis2020retrieval ; dong2025decoupling ; tu2025rbft ; su2025dynamic . By grounding the model in external knowledge, RAG directly addresses several fundamental limitations of LLMs, offering a robust mechanism to mitigate hallucinations ji2023survey ; su2024unsupervised ; su2024mitigating ; wang2025joint , facilitate knowledge updating fang2024alphaedit ; wang2024lekube ; wang2024knowledge ; wang2025decoupling , and enable effective domain adaptation yang2024tcm ; su2025judge ; su2024stard ; su2024legalaid .
The conventional approach to traditional RAG is built upon the "Retrieval-then-Read" paradigm borgeaud2022improving ; guu2020retrieval ; lewis2020retrieval . Within this framework, a user’s query triggers a search module for relevant documents within a large-scale external corpus. This retrieval step is carried out by either an external retriever zhai2008statistical ; su2023caseformer ; robertson2009probabilistic ; su2023wikiformer ; ma2023caseencoder ; fang2024scaling or a more sophisticated retrieval system su2023thuir2 ; salemi2024towards . Building upon this foundation, recent work has proposed more advanced RAG architectures to improve efficiency and effectiveness. For instance, Dynamic RAG jiang2022retrieval ; su2024dragin ; yao2024seakr moves beyond a single retrieval step by adaptively triggering the retriever during generation, specifically when the LLM is uncertain during the generation process. From another angle, GraphRAG edge2024local enhances the knowledge source by querying pre-constructed knowledge graphs instead of unstructured text, allowing it to retrieve interconnected facts and relationships. Furthermore, the Parametric RAG paradigm  su2025parametric ; tan2025dynamic ; fleshman2025lora alters the knowledge injection step by directly injecting retrieved knowledge into the LLM’s parameters.
The scientific survey generation task, which is the focus of our SurGE benchmark, presents a significant challenge for even these advanced RAG systems. Unlike typical question-answering, survey generation demands the synthesis of a large, diverse set of documents into a coherent, well-structured survey paper. Therefore, while RAG provides the foundational technology, our SurGE benchmark is specifically designed to push the boundaries of current models by rigorously evaluating their capabilities in large-scale multi-document synthesis and structured content creation.
7.2 Long-Form Text Generation and Evaluation
Long-form text generation is substantially more challenging than short-text generation due to its inherent requirements for sustained coherence and rich contextual understanding. Early approaches mainly used generative adversarial networks and reinforcement learning to conduct long-sequence generation (Guo2018, ). More recently, large language models have emerged as a strong tool for this task, offering advanced capabilities to handle long-text generation. For example, structured planning techniques and specialized inference mechanisms are proposed (Sloan2025, ; Jin2024, ) to generate consistent and high-quality clinical reports. Similarly, hierarchical planning frameworks have demonstrated that content control and multi-constraint instruction following can significantly enhance logical flow and topic coverage (hu2022, ; pham2024, ). Beyond medical or other task-specific applications, context-driven retrieval strategies, such as tree-structured retrieval, can support open-domain long-text generation by guiding the model through extensive knowledge sources (roy2024, ). The effective evaluation framework is vital for measuring the quality, factualness, and user-centric utility of the long-form text generation task. Traditional metrics, designed for shorter texts, often fail to capture the intricacies of longer outputs. Recent work has introduced task-focused benchmarks that emphasize user-oriented objectives, such as personalized writing or domain-specific content generation (kumar2024, ; salemi2025, ). In parallel, factuality assessment has attracted growing interest, with methods proposed to evaluate both verifiable and unverifiable claims. Metrics such as VERISCORE and FACTSCORE break down generated text into atomic facts, checking each for consistency against reliable sources (song2024, ; min2023, ). Beyond factual correctness, coherence and structural quality have been studied extensively. Benchmarks like LongGenBench and HelloBench underscore the importance of evaluating a model’s ability to maintain logical organization and clarity over extended passages (wu2025, ; que2024, ).
7.3 Survey Generation
In the domain of scientific writing, survey generation involves distilling extensive textual resources into a coherent and structured overview. Recent advances in AI-assisted systems have provided prompting-based approaches to expedite the drafting process while preserving content accuracy (gero2022, ; Kacena2024, ). One of the most commonly used approaches is retrieval-augmented generation, which combines large-scale knowledge retrieved from documents with language generation empowered by LLMs to yield factually comprehensive overviews (lewis2021, ). Retrieval-augmented generation is often initiated with dense retrieval methods based on dual-encoder architectures to identify highly relevant documents (karpukhin2020, ). Once these documents are retrieved, summarization techniques—spanning top-down, bottom-up, and graph-based ranking methods—play a pivotal role in producing concise yet faithful summaries (nayeem2024, ; pang2022, ; bleiweiss2023, ). Building on these retrieval and summarization-based methodologies, automated literature survey generation has garnered increasing attention (nlpcc2024, ; lai2024, ). However, existing techniques depend on limited ground truths and employ coarse evaluation metrics, resulting in oversimplified assessments of survey quality wang2024autosurvey . To address these challenges, we present a refined ground truth and a multi-dimensional evaluation framework that emphasizes both accuracy and structural coherence. By evaluating quality through multiple dimensions, our proposed framework advances the capabilities of automated survey generation, offering a more comprehensive and rigorous approach to summarizing scientific literature.
8 Conclusion
In this paper, we introduce SurGE, a benchmark for end-to-end scientific survey generation in computer science. We construct a high-quality dataset and develop a multi-dimensional evaluation framework that assesses the coverage, coherence, factual accuracy, and writing quality of automatically generated surveys. Extensive experiments with various baselines highlight the challenges in producing high-quality surveys. SurGE is expected to serve as a valuable resource for future research in information retrieval, fostering advancements in both retrieval methodologies and generative AI techniques. Several limitations guide future work, including: (1) The benchmark is limited in the domain of computer science but could be extended to broader domains. (2)  Integration with more advanced retrieval and generation methods could further enhance automated scientific survey generation for computer science.
Appendix A Details of the Academic Corpus
Table 3 provides an overview of the metadata schema used in our academic corpus. Each entry is structured to support efficient retrieval and interpretability.
Table 3: Fields and Descriptions for the Literature Knowledge Base.
Key Description
Title The title of the research paper.
Authors A list of contributing researchers.
Year The publication year of the paper.
Date The exact timestamp of the paper’s release.
Abstract The abstract of the paper.
Category The subject classification following the arXiv taxonomy.
doc_id A unique identifier assigned for reference and retrieval.
Appendix B Social Impact
Automated scientific survey writing brings two-sided social impacts. On the one hand, automated survey generation can save significant time and resources for researchers. This efficiency allows for quick data collection and analysis, enabling more rapid responses to a research domain. Automated survey generation could even provide innovation and lead to discoveries and insights with the help of LLMs. On the other hand, we have shown that the current method for automated survey generation is limited. Such a low-quality survey could poison the academic environment and bring ethical concerns, such as biased content generated by LLMs. Therefore, surveys generated by LLMs should be noted and published in different tracks. Such research direction should be carefully supervised by the collaboration of researchers.
Appendix C Selected Ground Truth Survey
The following three tables provide the complete list of the ground-truth surveys that comprise the SurGE benchmark. In line with our selection methodology outlined in Section 3, these surveys represent high-impact, peer-reviewed works. For each entry, we list its full title, publication year, primary arXiv category, and its citation count. The provided citation counts represent a snapshot taken from Google Scholar on May 10, 2025.
Survey Title Year Category Citation Count
A Survey on Edge Computing Systems and Tools 2019 cs.DC 384
A Survey on Graph-Based Deep Learning for Computational Histopathology 2021 cs.LG 141
A Survey of Uncertainty in Deep Neural Networks 2021 cs.LG 1547
A Survey on Explainability in Machine Reading Comprehension 2020 cs.CL 50
MAC Protocols for Terahertz Communication: A Comprehensive Survey 2019 cs.NI 154
Event Prediction in the Big Data Era: A Systematic Survey 2020 cs.AI 175
A Survey on Deep Neural Network Compression: Challenges, Overview, and Solutions 2020 cs.LG 142
Analysis of the hands in egocentric vision: A survey 2019 cs.CV 113
Neural Machine Translation for Low-Resource Languages: A Survey 2021 cs.CL 321
Physics-Guided Deep Learning for Dynamical Systems: A Survey 2021 cs.LG 115
A Survey on Bias and Fairness in Machine Learning 2019 cs.LG 6292
Generative Adversarial Networks for Spatio-Temporal Data: A Survey 2020 cs.LG 136
Ubiquitous Acoustic Sensing on Commodity IoT Devices: A Survey 2019 cs.SD 82
A Survey of Black-Box Adversarial Attacks on Computer Vision Models 2019 cs.LG 120
A Comprehensive Survey on Pretrained Foundation Models: A History from BERT to ChatGPT 2023 cs.AI 726
A Survey on Dynamic Network Embedding 2020 cs.SI 49
A Survey of Moving Target Defenses forNetwork Security 2019 cs.CR 295
A Survey on the Evolution of Stream Processing Systems 2020 cs.DC 111
Deep Gait Recognition: A Survey 2021 cs.CV 264
Transformers in Vision: A Survey 2021 cs.CV 3276
Image Classification with Deep Learning in the Presence of Noisy Labels: A Survey 2019 cs.LG 433
Change Detection and Notification of Web Pages: A Survey 2019 cs.IR 26
A Survey on Deep Learning-based Architecturesfor Semantic Segmentation on 2D images 2019 cs.CV 276
A Survey on Tiering and Caching in High-Performance Storage Systems 2019 cs.AR 28
Multimodal Learning with Transformers: A Survey 2022 cs.CV 766
Attention, please! A survey of Neural Attention Models in Deep Learning 2021 cs.LG 251
Explanation-Based Human Debugging of NLP Models: A Survey 2021 cs.CL 80
Federated Learning in Mobile Edge Networks: A Comprehensive Survey 2019 cs.NI 2488
Deep Learning for Image Super-resolution:A Survey 2019 cs.CV 2036
Deep Learning for Weakly-Supervised Object Detection and Object Localization: A Survey 2021 cs.CV 25
Survey of Transient Execution Attacks 2020 cs.CR 23
A Survey of Syntactic-Semantic Parsing Based on Constituent and Dependency Structures 2020 cs.CL 47
Efficient Deep Learning: A Survey on Making Deep Learning Models Smaller, Faster, and Better 2021 cs.LG 519
From Statistical Relational to Neurosymbolic Artificial Intelligence: a Survey. 2021 cs.AI 71
Symbolic Logic meets Machine Learning: A Brief Survey in Infinite Domains 2020 cs.AI 50
Context Dependent Semantic Parsing: A Survey 2020 cs.CL 21
A survey of active learning algorithms for supervised remote sensing image classification 2021 cs.CV 651
Generate FAIR Literature Surveys with Scholarly Knowledge Graphs 2020 cs.DL 53
A Survey of Deep Learning for Data Caching in Edge Network 2020 cs.NI 39
Logic Locking at the Frontiers of Machine Learning: A Survey on Developments and Opportunities 2021 cs.CR 29
Affective Computing for Large-Scale Heterogeneous Multimedia Data: A Survey 2019 cs.MM 95
Weakly Supervised Object Localization and Detection: A Survey 2021 cs.CV 348
Compression of Deep Learning Models for Text: A Survey 2020 cs.CL 142
Computer Vision with Deep Learning for Plant Phenotyping in Agriculture: A Survey 2020 cs.CV 85
A Survey on Evolutionary Neural Architecture Search 2020 cs.NE 641
Towards Efficient SynchronousFederated Training: A Survey onSystem Optimization Strategies 2021 cs.DC 34
From Distributed Machine Learning to Federated Learning: A Survey 2021 cs.DC 342
Low-Light Image and Video Enhancement Using Deep Learning: A Survey 2021 cs.CV 522
A Survey of Coded Distributed Computing 2020 cs.DC 28
A Systematic Survey of Regularization and Normalization in GANs 2020 cs.LG 56
Deep Learning for Deepfakes Creation and Detection: A Survey 2019 cs.CV 614
Proximity Perception in Human-Centered Robotics: A Survey on Sensing Systems and Applications 2021 cs.RO 127
A Survey of Transformers 2021 cs.LG 1641
How should my chatbot interact? A survey on social characteristics in human-chatbot interaction design 2019 cs.HC 676
[…]
A Survey of Deep Reinforcement Learning in Recommender Systems: A Systematic Review 2021 cs.IR 73
Learning from Noisy Labels with Deep Neural Networks: A Survey 2020 cs.LG 1380
A Survey on Split Manufacturing: Attacks, Defenses, and Challenges 2020 cs.CR 59
A Survey of Active Learning for Text Classification using Deep Neural Networks 2020 cs.CL 145
A Survey of Knowledge Tracing: Models, Variants, and Applications 2021 cs.CY 51
Dynamic Neural Networks: A Survey 2021 cs.CV 849
Arms Race in Adversarial Malware Detection: A Survey 2020 cs.CR 81
A Comprehensive Survey on Graph Anomaly Detection with Deep Learning 2021 cs.LG 775
Deep Learning for Instance Retrieval: A Survey 2021 cs.CV 292
Deep Learning for Vision-based Prediction: A Survey 2020 cs.CV 59
A Survey of Machine Learning for Computer Architecture and Systems 2021 cs.LG 91
Inspect, Understand, Overcome:A Survey of Practical Methodsfor AI Safety 2021 cs.LG 78
Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394
A Survey on Recent Approaches for Natural Language Processing in Low-Resource Scenarios 2020 cs.CL 390
DL-Traff: Survey and Benchmark of Deep Learning Models for Urban Traffic Prediction 2021 cs.LG 180
Survey Title Year Category Citation Count
A survey of face recognition techniques under occlusion 2020 cs.CV 165
Blockchain for 5G and Beyond Networks: A State of the Art Survey 2019 cs.NI 440
A Survey of Knowledge-Enhanced Text Generation 2020 cs.CL 344
Hardware Acceleration of Sparse and Irregular Tensor Computations of ML Models: A Survey 2020 cs.AR 112
Graph Learning for Combinatorial Optimization: A Survey of State-of-the-Art 2020 cs.LG 113
Deep Gaussian Processes: A Survey 2021 cs.LG 39
Graph Learning: A Survey 2021 cs.LG 515
Harnessing the Power of LLMs in Practice: A Survey on ChatGPT and Beyond 2023 cs.CL 981
A Survey on In-context Learning 2022 cs.CL 1788
Centrality Measures in Complex Networks: A Survey 2020 cs.SI 139
A Survey on Adversarial Recommender Systems 2020 cs.IR 244
Towards a Robust Deep Neural Network in Texts: A Survey 2019 cs.CL 90
A Survey of Deep Active Learning 2020 cs.LG 1588
Translation Quality Assessment: A Brief Survey on Manual and Automatic Methods 2021 cs.CL 56
Deep learning for scene recognition from visual data: a survey 2020 cs.CV 23
A Survey of State-of-the-Art on Blockchains: Theories, Modelings, and Tools 2020 cs.DC 170
Computation Offloading and Content Caching Delivery in Vehicular Edge Computing: A Survey 2019 cs.NI 101
Recent Advances in Deep Learning Based Dialogue Systems: A Systematic Survey 2021 cs.CL 322
Explainable reinforcement learning for broad-XAI: a conceptual framework and survey 2021 cs.AI 82
End-to-End Constrained Optimization Learning: A Survey 2021 cs.LG 258
Machine Learning in Generation, Detection, and Mitigation of Cyberattacks in Smart Grid: A Survey 2020 cs.CR 31
Pervasive AI for IoT applications: A Survey on Resource-efficient Distributed Artificial Intelligence 2021 cs.DC 153
A Survey of Deep Learning Approaches for OCR and Document Understanding 2020 cs.CL 84
Taxonomy of Machine Learning Safety: A Survey and Primer 2021 cs.LG 81
Neuron-level Interpretation of Deep NLP Models: A Survey 2021 cs.CL 98
Using Deep Learning to Solve Computer Security Challenges: A Survey 2019 cs.CR 59
Efficient Transformers: A Survey 2020 cs.LG 1500
A Survey of Label-noise Representation Learning: Past, Present and Future 2020 cs.LG 208
A Survey of Constrained Gaussian Process Regression: Approaches and Implementation Challenges 2020 cs.LG 182
Neural Networks for Entity Matching: A Survey 2020 cs.DB 157
A Survey of Orthographic Information in Machine Translation 2020 cs.CL 37
A Survey of Quantum Theory Inspired Approaches to Information Retrieval 2020 cs.IR 42
A Survey of Deep Meta-Learning 2020 cs.LG 459
A Survey on Deep Learning Techniques for Video Anomaly Detection 2020 cs.CV 46
A Survey on Interactive Reinforcement Learning 2021 cs.HC 128
Applications of Auction and Mechanism Design in Edge Computing: A Survey 2021 cs.GT 76
Abduction and Argumentation for Explainable Machine Learning: A Position Survey 2020 cs.AI 19
A Survey of Exploration Methods in Reinforcement Learning 2021 cs.LG 133
Reinforcement Learning based Recommender Systems: A Survey 2021 cs.IR 597
A Survey of Data Augmentation Approaches for NLP 2021 cs.CL 1008
A Survey on Self-supervised Pre-training for Sequential Transfer Learning in Neural Networks 2020 cs.LG 77
A Survey on Theorem Provers in Formal Methods 2019 cs.SE 45
A Survey on Heterogeneous Graph Embedding 2020 cs.SI 452
Asynchronous Federated Learning on Heterogeneous Devices: A Survey 2021 cs.DC 326
A Survey on Data-driven Software Vulnerability Assessment and Prioritization 2021 cs.SE 94
Time Series Data Imputation: A Survey on Deep Learning Approaches 2020 cs.LG 96
A Survey on Low-Resource Neural Machine Translation 2021 cs.CL 59
Benchmark and Survey of Automated Machine Learning Frameworks 2019 cs.LG 523
A Survey on Deep Learning Technique for Video Segmentation 2021 cs.CV 204
Universal Adversarial Perturbations: A Survey 2020 cs.CV 66
A Survey of Machine Learning Methods and Challenges for Windows Malware Classification 2020 cs.CR 73
Serverless Computing: A Survey of Opportunities, Challenges, and Applications 2019 cs.NI 295
A Survey on Subgraph Counting: Concepts, Algorithms and Applications to Network Motifs and Graphlets 2019 cs.DS 218
Survey of Attacks and Defenses on Edge-Deployed Neural Networks 2019 cs.CR 51
Reinforcement learning with human advice: a survey. 2020 cs.AI 86
Domain Generalization: A Survey 2021 cs.LG 1487
Deep Learning for 3D Point Clouds: A Survey 2019 cs.CV 2360
Survey on Causal-based Machine Learning Fairness Notions 2020 cs.LG 109
Towards a Survey on Static and Dynamic Hypergraph Visualizations 2021 cs.HC 38
Graph-based Deep Learning for Communication Networks: A Survey 2021 cs.NI 245
Opportunities and Challenges in Explainable Artificial Intelligence (XAI): A Survey 2020 cs.CV 1062
Domain Adaptation and Multi-Domain Adaptation for Neural Machine Translation: A Survey 2021 cs.CL 108
Survey Title Year Category Citation Count
What Can Knowledge Bring to Machine Learning? – A Survey of Low-shot Learning for Structured Data 2021 cs.LG 28
A Survey of Knowledge Graph Embedding and Their Applications 2021 cs.IR 68
Deep Graph Generators: A Survey 2020 cs.LG 75
Post-hoc Interpretability for Neural NLP: A Survey 2021 cs.CL 309
A Survey on Machine Reading Comprehension: Tasks, Evaluation Metrics and Benchmark Datasets 2020 cs.CL 129
A survey on bias in visual datasets 2021 cs.CV 153
A Comprehensive Survey on Graph Neural Networks 2019 cs.LG 12047
A survey of security and privacy issues in the Internet of Things from
the layered context
2019 cs.NI 144
Adversarial Machine Learning in Image Classification: A Survey Towards the Defender’s Perspective 2020 cs.CV 217
Survey of Machine Learning Accelerators 2020 cs.DC 230
Online Fair Division: A Survey 2019 cs.AI 55
A Comprehensive Survey of Scene Graphs: Generation and Application 2021 cs.CV 354
A Survey of Numerical Methods Utilizing Mixed Precision Arithmetic 2020 cs.MS 216
Image Data Augmentation for Deep Learning: A Survey 2022 cs.CV 445
Neural Unsupervised Domain Adaptation in NLP—A Survey 2020 cs.CL 357
A Survey of Evaluation Metrics Used for NLG Systems 2020 cs.CL 305
Deep Learning for Micro-expression Recognition: A Survey 2021 cs.CV 99
A Survey of Human-in-the-loop for Machine Learning 2021 cs.LG 735
Deep Fake Detection : Survey of Facial Manipulation Detection Solutions 2021 cs.CV 36
Survey of Hallucination in Natural Language Generation 2022 cs.CL 3976
The Rise and Potential of Large Language Model Based Agents: A Survey 2023 cs.AI 971
Generalized Out-of-Distribution Detection: A Survey 2021 cs.CV 1158
A survey on practical adversarial examples for malware classifiers 2020 cs.CR 23
One-Class Classification: A Survey 2021 cs.CV 196
Survey of Low-Resource Machine Translation 2021 cs.CL 197
A Survey on Visual Transformer 2020 cs.CV 405
Automatic Speech Recognition And Limited Vocabulary: A Survey 2021 cs.AI 61
A Comprehensive Survey on Community Detection with Deep Learning 2021 cs.SI 487
A Survey of Distributed Consensus Protocols for Blockchain Networks 2019 cs.CR 1101
Spatiotemporal Data Mining: A Survey on Challenges and Open Problems 2021 cs.LG 148
RGB-D Salient Object Detection: A Survey 2020 cs.CV 305
A Survey on Embedding Dynamic Graphs 2021 cs.LG 177
Transformers in Time Series: A Survey 2022 cs.LG 1152
Deep Learning for 3D Point Cloud Understanding: A Survey 2020 cs.CV 47
AMMU : A Survey of Transformer-based Biomedical Pretrained Language Models 2021 cs.CL 170
Instruction Tuning for Large Language Models: A Survey 2023 cs.CL 919
A Short Survey of Pre-trained Language Models for Conversational AI-A New Age in NLP 2021 cs.CL 94
A Survey on Evaluation of Large Language Models 2023 cs.CL 2885
D’ya like DAGs? A Survey on Structure Learning and Causal Discovery 2021 cs.LG 364
Machine Learning Testing: Survey, Landscapes and Horizons 2019 cs.LG 1059
Deep Learning-based Spacecraft Relative Navigation Methods: A Survey 2021 cs.RO 100
A Survey on Deep Learning Event Extraction: Approaches and Applications 2021 cs.CL 107
Graph Neural Networks for Natural Language Processing: A Survey 2021 cs.CL 396
Stabilizing Generative Adversarial Networks: A Survey 2019 cs.LG 149
A Survey of Knowledge-based Sequential Decision Making under Uncertainty 2020 cs.AI 26
Automated Machine Learning on Graphs: A Survey 2021 cs.LG 97
Active Divergence with Generative Deep Learning - A Survey and Taxonomy 2021 cs.LG 23
[…]
A Survey on Multi-modal Summarization 2021 cs.CL 73
Generative adversarial networks in time series: A survey and taxonomy 2021 cs.LG 105
Towards Reasoning in Large Language Models: A Survey 2022 cs.CL 841
Survey on reinforcement learning for language processing 2021 cs.CL 180
Creativity and Machine Learning: A Survey 2021 cs.LG 56
A Survey of Deep Reinforcement Learning Algorithms for Motion Planning and Control of Autonomous Vehicles 2021 cs.RO 92
Deep Reinforcement Learning in Computer Vision: A Comprehensive Survey 2021 cs.CV 238
Domain Name System Security and Privacy: A Contemporary Survey 2020 cs.NI 69
Resource Allocation and Service Provisioning in Multi-Agent Cloud Robotics: A Comprehensive Survey 2021 cs.RO 145
Generalizing to Unseen Domains: A Survey on Domain Generalization 2021 cs.LG 1311
Deep Graph Similarity Learning: A Survey 2019 cs.LG 103
Word Embeddings: A Survey 2019 cs.CL 449
Augmented Language Models: a Survey 2023 cs.CL 544
Software Engineering for AI-Based Systems: A Survey 2021 cs.SE 317
Pre-train, Prompt, and Predict: A Systematic Survey of Prompting Methods in Natural Language Processing 2021 cs.CL 5613
A Survey on Knowledge Graph Embeddings with Literals: Which model links better Literal-ly? 2019 cs.AI 97
A comprehensive taxonomy for explainable artificial intelligence 2021 cs.LG 276
Federated Learning Meets Natural Language Processing: A Survey 2021 cs.CL 103
Fairness in Machine Learning: A Survey 2020 cs.LG 1163
Deep Learning for Text Style Transfer: A Survey 2020 cs.CL 318
Retrieval-Augmented Generation for Large Language Models: A Survey 2023 cs.CL 1951
Transformers in Medical Imaging: A Survey 2022 eess.IV 889
A Survey of Privacy Attacks in Machine Learning 2020 cs.CR 368
A Survey on Deep Reinforcement Learning for Data Processing and Analytics 2021 cs.LG 44
A Survey on Dialogue Summarization: Recent Advances and New Frontiers 2021 cs.CL 111

### v2

@@
-Benchmarking Computer Science Survey Generation
+SurGE: A Benchmark and Evaluation Framework for Scientific Survey Generation
 Abstract
-Scientific survey articles play a vital role in summarizing research progress, yet their manual creation is becoming increasingly infeasible due to the rapid growth of academic literature.
-While large language models (LLMs) offer promising capabilities for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols.
-To address this gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for evaluating scientific survey generation in the computer science domain.
-SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers that serves as the retrieval pool.
-In addition, we propose an automated evaluation framework that measures generated surveys across four dimensions: information coverage, referencing accuracy, structural organization, and content quality.
-Our evaluation of diverse LLM-based approaches shows that survey generation remains highly challenging, even for advanced self-reflection frameworks.
-These findings highlight the complexity of the task and the necessity for continued research [^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
+The rapid growth of academic literature makes the manual creation of scientific surveys increasingly infeasible.
+While large language models show promise for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols.
+To bridge this critical gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for scientific survey generation in computer science.
+SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers.
+In addition, we propose an automated evaluation framework that measures the quality of generated surveys across four dimensions: comprehensiveness, citation accuracy, structural organization, and content quality.
+Our evaluation of diverse LLM-based methods demonstrates a significant performance gap, revealing that even advanced agentic frameworks struggle with the complexities of survey generation and highlighting the need for future research in this area.
+[^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
 1 Introduction
-Large language models (LLMs) have shown impressive capabilities in scientific text generation wang2024autosurvey ; gao2024evaluating ; liang2025surveyx , supporting various stages of academic writing from outlining to full-content generation.
-Among these applications, automated scientific survey generation stands out as a particularly valuable task.
-Scientific survey articles (also known as survey papers or literature reviews) synthesize knowledge from a broad collection of academic papers into a structured overview of a research area keele2007guidelines ; pautasso2013ten ; boote2005scholars .
-Traditionally, writing such surveys is highly labor-intensive: an expert in this field must gather dozens or even hundreds of relevant papers, understand their contributions, identify important themes and trends, and finally craft a well-structured survey paper.
-This labor-intensive process has become increasingly challenging due to the explosive growth of academic literature in recent years.
-For example, arXiv now receives over a thousand new computer science papers each day, and the annual volume of submissions has more than doubled from 2019 to 2024 liang2025surveyx .
-Keeping pace with this scale of information is infeasible for individual researchers, prompting growing interest in automatic survey writing.
-Despite the success of LLMs in various generation tasks, generating structured and comprehensive scientific surveys has received relatively little attention.
-A key reason is the lack of standardized benchmarks and evaluation protocols tailored to the unique demands of this task.
-Unlike general-purpose generation tasks, scientific survey writing remains difficult to evaluate in a quantitative and standardized way, due to the lack of benchmarks that capture its unique requirements: synthesizing multiple sources, ensuring citation accuracy, and maintaining structural coherence.
-As a result, current systems are often assessed through manual expert reviews nlpcc2024 , which are costly, time-consuming, and hinder reproducibility and systematic progress.
-To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark designed to facilitate reproducible research in scientific survey generation.
-We formalize the task as a two-stage process: given a topic description in the computer science domain (e.g., Machine Learning for Information Retrieval), the system should first collect a set of relevant papers from a large academic corpus and then synthesize a well-structured survey from them.
-To support this formulation, SurGE provides the necessary resources and a standardized evaluation framework for both the retrieval and generation stages.
-Each test instance in SurGE comprises a research topic and its corresponding ground-truth survey, which serves as a high-quality reference for evaluation.
-These ground-truth surveys are carefully selected from peer-reviewed, high-impact survey papers, based on citation statistics and topical relevance.
-SurGE dataset consists of two main components: (1) a set of test instances, where each instance includes a topic description, its ground-truth survey, and the full list of cited references in that survey; (2) a large-scale academic corpus containing over one million computer science papers, which serves as the retrieval pool for relevant document selection and coverage analysis.
-Beyond dataset construction, we also propose an automated evaluation framework that enables reproducible and fine-grained assessment of generated surveys.
-This framework scores outputs across four key dimensions: information collection, referencing accuracy, structural organization, and content quality.
-Grounded in established academic writing guidelines webster2002analyzing ; boote2005scholars ; keele2007guidelines ; pautasso2013ten , this framework enables automatic, scalable evaluation without the need for human annotation, making it well-suited for benchmarking and driving systematic progress in this emerging task.
+The volume of scientific literature has been expanding at an unprecedented rate in recent years.
+For instance, academic archives like arXiv now receive over a thousand new computer science papers daily, more than doubled between 2019 and 2024 (Liang et al., 2025).
+This rapid growth in publications has made the manual creation of comprehensive survey papers increasingly impractical, as the manual collection and synthesis of large volumes of relevant papers is both labor-intensive and time-consuming.
+Faced with this challenge, there is a growing need for automated systems that can generate survey papers effectively.
+While recent advancements in LLM-based agents (Yao et al., 2023; Wang et al., 2024c) offer significant promise for automating this task, their full potential is severely hindered by the lack of reliable, scalable, and standardized evaluation benchmarks.
+Currently, the evaluation of automatically generated surveys largely depends on human-expert reviews, which limit the reproducibility and objectivity of assessments (Tian et al., 2024a).
+Consequently, progress in this field is difficult to quantify, and the relative merits of different methods remain challenging to compare without standardized, multi-faceted evaluation benchmarks.
+To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark that establishes a reproducible standard for the survey generation task, which we formalize as a two-stage process: (1) retrieving relevant papers for a given topic from a large corpus, and (2) synthesizing them into a well-structured survey.
+To support standardized evaluation of both stages, SurGE provides a comprehensive dataset consisting of two key components.
+The first is a large-scale academic corpus containing over one million computer science papers to be used as the search and retrieval pool.
+The second is a collection of test instances, where each instance comprises a research topic (e.g., ”Machine Learning for Information Retrieval”) and its corresponding ground-truth survey.
+Each ground-truth survey is a high-impact, peer-reviewed paper, chosen based on strict citation criteria and subsequently validated for its quality by expert annotators (see Section 3.1), and is provided along with its full list of cited references.
+Beyond dataset construction, another key contribution of this work is a fully automated evaluation framework that enables reproducible and multi-faceted assessment of generated surveys.
+Grounded in established principles for high-quality survey writing (Webster & Watson, 2002; Boote & Beile, 2005; Keele et al., 2007; Pautasso, 2013), our proposed evaluation framework quantifies a survey’s quality across four key dimensions: comprehensiveness, citation accuracy, structural organization, and content quality.
+This automated approach eliminates the reliance on subjective and time-consuming human evaluation, thus establishing a scalable and standardized methodology to guide and measure progress in survey generation.
 To validate the utility of SurGE, we evaluate a range of LLM-based baseline systems on the benchmark.
 These include standard retrieval-augmented generation approaches that first retrieve relevant documents and then generate a survey, as well as more advanced pipelines that incorporate explicit planning (e.g., outline generation) and iterative refinement of the draft.
-The results reveal that even the state-of-the-art systems struggle with the survey generation task, highlighting its difficulty.
-For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant references.
-These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research.
-We hope that SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
+The results reveal that even the state-of-the-art systems struggle with the survey generation task.
+For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant citations.
+These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research in this area.
+We believe SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
 In summary, our contributions are threefold:
   1. 1.
-    We introduce SurGE, the first comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
+    We introduce SurGE, a comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
   2. 2.
-    We propose a fully automated evaluation framework that assesses survey quality across four crucial dimensions: information coverage, citation accuracy, structure, and content.
+    We propose an automated evaluation framework that assesses survey quality across four crucial dimensions: comprehensiveness, citation accuracy, structure, and content.
   3. 3.
@@
 The process consists of:
-  • Document Collection: A retrieval module or complex retrieval system identifies a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
+  • Document Collection: A retrieval module or complex agentic retrieval system collects a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
   • Survey Generation: A generative model composes a well-structured survey $S$ based on the topic $t$ and the retrieved document set $\mathcal{R}_{t}$, including proper citations and a reference list.
-The output survey $S$ is expected to (i) follow a logical hierarchical organization, (ii) cover key techniques and trends related to the topic, and (iii) provide accurate citations in the appropriate places using a consistent and correct format.
+To enable standardized benchmarking of the survey generation task, SurGE offers a comprehensive framework that formalizes the process.
+The benchmark includes 205 expert-selected topics, each paired with a topic description $t$ and a ground-truth survey for evaluation.
+SurGE also provides a large-scale academic corpus $D$ containing over one million papers, which supports the document retrieval stage.
+The generated surveys are then evaluated through an automated system, which quantitatively assesses their quality across four key dimensions.
+Details of the dataset construction process are outlined in Section 3, while the evaluation framework is explained in Section 4.
 3 Dataset Construction
-3.1 Ground-Truth Survey Collection and Annotation
+3.1 Ground-Truth Survey Collection and Expert Annotation
 To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature.
 Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews.
-To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact bornmann2008citation ; and (iii) the publish date must be between 2020 and 2024.
-Following the initial filtering process, we further refined the SurGE dataset through expert annotation to ensure the inclusion of only high-quality survey documents.
-This stage aimed to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers.
+To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact (Bornmann & Daniel, 2008); and (iii) the publication date must be between 2020 and 2024.
+Following the initial filtering process, we further refine the SurGE dataset through expert annotation.
+This process aims to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers.
 To this end, we recruited a team of four computer science Ph.D.
@@
 Annotators could typically complete the task within 8–9 minutes per paper on average.
-Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) reference quality, which examines the relevance, diversity, and traceability of cited works.
+Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) citation quality, which examines the relevance, diversity, and traceability of cited works.
 Each annotator labeled the document as either “usable” or “not usable.” A survey paper was included in the final dataset only if both annotators independently marked it as usable.
@@
 After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers.
-The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix C.
-3.2 Reference Extraction and Processing
-[…]
-The core reason for extracting the full reference list from each ground-truth survey is to create a gold standard for evaluating the information collection stage.
-We operate on the premise that the citations in an expert-written survey represent a curated collection of the field’s most foundational and relevant literature.
-By using this set as the ground truth, we can quantitatively measure the coverage and recall of a system-generated survey’s references, providing a clear metric for its performance.
+The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix F.
+For each of the 205 selected ground-truth surveys, we provide its complete reference list, which serves as the gold standard for evaluating the comprehensiveness of a generated survey.
+The detailed methodology for extracting and processing these references is described in Appendix A.
 3.2 Academic Corpus Construction
@@
 Our corpus is built entirely from scholarly metadata obtained from the arXiv repository.
-To ensure adherence to ethical and legal standards, we exclusively collected metadata and did not include full-text PDFs, a practice permitted by arXiv’s Terms of Use, which designates metadata as public domain under the CC0 license arxiv_tou .
+To ensure adherence to ethical and legal standards, we exclusively collected metadata and did not include full-text PDFs, a practice permitted by arXiv’s Terms of Use, which designates metadata as public domain under the CC0 license (arx, 2025b).
 The corpus was constructed through a two-stage process.
@@
 We do not host or redistribute arXiv PDFs or source files.
-This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication arxiv_tou .
-This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata arxiv_submission .
+This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication (arx, 2025b).
+This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata (arx, 2025a).
 We have prioritized transparency, reproducibility, and ethical considerations throughout dataset construction.
@@
 Regular updates will ensure continued relevance and alignment with evolving research trends and standards.
+Figure 1: Overview of the SurGE Benchmark. (a) Summary statistics of the curated survey dataset and its associated retrieval corpus. (b) Metadata of the pre-processed survey dataset used in SurGE.
 (a) Basic Statistics of the SurGE Dataset and Corpus
@@
 All_Cites List of document IDs cited in the survey.
-Figure 1: Overview of the SurGE Benchmark. (a) Summary statistics of the curated survey dataset and its associated retrieval corpus. (b) Metadata of the pre-processed survey dataset used in SurGE.
 4 Evaluation Framework
-Developing a robust evaluation framework for automated scientific survey generation involves identifying critical dimensions that reflect a survey’s overall accuracy, comprehensiveness, and quality.
-Through discussions with academic experts, peer reviewers, and experienced survey authors, we established four essential evaluation dimensions:Information Collection Quality, Referencing Accuracy, Structural Quality, and Content Quality.
-Each generated survey is systematically compared with an expert-written Ground Truth (GT) survey, ensuring alignment with high academic standards.
-In the following subsections, we deeply analyze each evaluation criterion and subsequently define the corresponding quantitative metrics.
-4.1 Information Collection Quality
-A fundamental requirement for scientific surveys is the comprehensive collection of relevant literature, covering key works that accurately represent the research field.
-Omissions or insufficient coverage can severely compromise a survey’s credibility, causing readers to question its thoroughness.
-Therefore, this aspect is quantitatively assessed through the Coverage Score, measuring the extent to which a generated survey includes key literature identified by expert consensus in the GT survey.
-Metric Definition.
-The coverage score $\mathcal{C}$ is defined as the recall of references included in the generated survey relative to the GT survey:
-$\mathcal{C}=\frac{|{R}_{GT}\cap{R}_{G}|}{|{R}_{GT}|},$ (1)
-where ${R}_{GT}$ denotes references in the GT survey, and ${R}_{G}$ denotes references in the generated survey. A higher coverage score indicates more comprehensive inclusion of essential literature.
-While the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited, peer-reviewed publications.
-We therefore interpret the Coverage score not as a measure of absolute completeness, but as a pragmatic metric for evaluating a system’s ability to identify the central body of work validated by the research community.
-4.2 Referencing Accuracy
-Referencing Accuracy evaluates how well the references cited in a generated survey align with its topic and content.
-This assessment is crucial for ensuring that the survey is well-supported and that citations are contextually appropriate.
-We propose a hierarchical evaluation framework that assesses referencing quality from three perspectives:
-  • Document-level: Determines if a cited paper is thematically relevant to the survey’s overall topic.
-  • Section-level: Evaluates if a reference is placed within a semantically appropriate section of the survey.
-  • Sentence-level: Assesses if a reference directly supports the specific claim made in the sentence where it is cited.
-4.2.1 Automated Relevance Judgment.
-To automate this evaluation, we employ a Natural Language Inference (NLI) model (cross-encoder/nli-deberta-v3-base) to automatically assess relevance.
-For each reference, we form a premise-hypothesis pair and use the NLI model to predict the relationship between them.
-The premise is constructed from the reference’s metadata, while the hypothesis is contextualized by its placement in the generated survey.
-Specifically, a reference is judged based on the NLI model’s output scores for “Entailment,” “Neutral,” and “Contradiction”:
-  • It is deemed relevant if “Entailment” has the highest score.
-  • It is deemed partially relevant if “Neutral” is highest and “Entailment” is the second-highest.
-  • Otherwise, it is deemed irrelevant.
-Since the survey generation process is grounded in our academic corpus, any reference cited by the model that is not found within this corpus is considered a hallucination.
-Such references are therefore automatically marked as irrelevant across all three hierarchical levels.
-4.2.2 Hierarchical Evaluation Metrics.
-1.
-Document-level Accuracy ($R_{d}$).
-This metric assesses if each reference is relevant to the survey’s main topic.
-For a cited reference $r$, we first apply a heuristic: if $r$ is also cited in the ground-truth survey, it is automatically marked as relevant.
-For all other references, we use the NLI model with the following pair:
-  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
-  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract could be cited in the paper: ‘{Survey Title}’.”
-2.
-Section-level Accuracy ($R_{s}$).
-This metric evaluates the contextual appropriateness of a citation within a specific section.
-We assess each unique (reference, section) pair.
-If a citation marker in the text points to a non-existent entry in the bibliography, the pair is marked as irrelevant.
-Otherwise, we query the NLI model:
-  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
-  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract is relevant to the section: ‘{Section Title}’.”
-3.
-Sentence-level Accuracy ($R_{t}$).
-For the most fine-grained evaluation, this metric measures if a reference directly supports the claim in its citing sentence.
-We assess each unique (reference, sentence) pair using the NLI model:
-  • Premise: “There is a paper. Title: ‘{Reference Title}’. Abstract: {Reference Abstract}.”
-  • Hypothesis: “The paper titled ‘{Reference Title}’ with the given abstract could be cited in the sentence: ‘{Sentence Text}’.”
-4.2.3 Metric Calculation.
-Based on the automated judgments, we define indicator functions for each reference $r$ at a given level: $\mathbb{I}_{\text{rel}}(r)=1$ if $r$ is relevant, and 0 otherwise; $\mathbb{I}_{\text{half}}(r)=1$ if $r$ is partially relevant, and 0 otherwise.
-The final accuracy score for each level $x\in\{d,s,t\}$ is calculated as:
-$R_{x}=\frac{\left|\{r\in\mathcal{R}_{x}\mid r\text{ is relevant}\}\right|+0.5\cdot\left|\{r\in\mathcal{R}_{x}\mid r\text{ is partially relevant}\}\right|}{\left|\mathcal{R}_{x}\right|},$ (2)
-where $\mathcal{R}_{d}$ is the set of all unique references in the generated survey, while $\mathcal{R}_{s}$ and $\mathcal{R}_{t}$ are the sets of all unique (reference, section) and (reference, sentence) citation instances, respectively.
+To comprehensively evaluate the quality of automatically generated scientific surveys, we propose a multi-faceted evaluation framework.
+This framework assesses survey quality across four crucial dimensions: Comprehensiveness, Citation Accuracy, Structural Quality, and Content Quality.
+Each generated survey is evaluated against an expert-written Ground Truth (GT) survey.
+The following subsections define the quantitative metrics for each dimension in detail.
+4.1 Comprehensiveness
+The comprehensiveness of a scientific survey is a critical quality factor, as the omission of key publications can undermine its credibility and value.
+To quantify this aspect, we evaluate the Recall of a generated survey’s references against the ground-truth reference lists.
+Formally, let ${R}_{GT}$ be the set of references in an expert-written GT survey and ${R}_{G}$ be the set of references in our generated survey.
+Recall $\mathcal{R}$ is defined as:
+$\mathcal{R}=\frac{|{R}_{GT}\cap{R}_{G}|}{|{R}_{GT}|},$ (1)
+While the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited, peer-reviewed publications (detailed in §3.1).
+We therefore interpret this metric not as a measure of absolute completeness, but as a pragmatic metric for evaluating a system’s ability to identify the central body of work validated by the research community.
+4.2 Citation Accuracy
+Citation accuracy is another critical aspect of a high-quality survey.
+Each citation in a survey should be thematically relevant to the overall topic of the survey, and it must be contextually appropriate in terms of both the section and sentence in which it appears.
+To evaluate this aspect, we introduce the metric of Citation Accuracy, which evaluates each citation in the survey across three levels.
+First, at the document level, we assess whether a cited paper is thematically relevant to the overall topic of the survey.
+Second, at the section level, we evaluate whether a citation is placed in a semantically appropriate section of the survey.
+Finally, at the sentence level, we verify whether a citation supports the specific claim made in the sentence where it is cited.
+To automate this evaluation, we employ a Natural Language Inference (NLI) model (nli-deberta-v3-base [^3] https://huggingface.co/cross-encoder/nli-deberta-v3-base) to assess the relevance of each citation.
+An NLI model is designed to determine the logical relationship between two text snippets: a premise and a hypothesis.
+The model then predicts the relationship between these two components, providing probabilities for the following labels: Entailment, Neutral, Contradiction.
+Due to its ability to capture semantic relationships, NLI has become a standard method for evaluating the factual consistency and relevance of text generated by LLMs, rendering it an ideal tool for our task.
+For our specific evaluation, we implement this three-level check by framing it as a series of NLI tasks.
+For each citation $r$ (with title $T_{r}$ and abstract $A_{r}$) within the generated survey $S$ (with title $T_{S}$), we construct a set of premise-hypothesis pairs.
+The premise is consistently formulated using the content of the cited paper, providing the evidentiary basis for the claim.
+The hypothesis is specifically tailored to assert relevance at each of the three levels (document, section, and sentence).
+This formulation is structured as follows:
+NLI Task Formulation Premise (Consistent for all levels): There is a paper.
+Title: $T_{r}$”.
+Abstract: $A_{r}$.
+Hypotheses (Tailored for each level of granularity): Document-level: The paper titled $T_{r}$” with the given abstract is thematically relevant to the survey titled: $T_{S}$”.
+Section-level: The paper titled $T_{r}$” with the given abstract is relevant to the section: “Section Title”.
+Sentence-level: The paper titled $T_{r}$” with the given abstract supports the claim made in the sentence: “Sentence Text”.
+The score for each citation unit is calculated via a multi-step process.
+Let $R$ denote the set of all citation instances in the generated survey.
+For each citation $r\in R$, we compute a score at each of the three levels: document ($y_{d}(r)$), section ($y_{s}(r)$), and sentence ($y_{t}(r)$).
+The calculation proceeds as follows.
+First, we resolve two special cases without querying the NLI model.
+Any citation $r$ not found in our academic corpus is classified as a hallucination and assigned a score of $y_{x}(r)=0$ at all levels $x\in\{d,s,t\}$.
+Conversely, any citation $r$ that is present in the ground-truth survey’s bibliography is assigned a document-level score of $y_{d}(r)=1$.
+For all other cases, the relevance score $y_{x}(r)$ is determined by the NLI model’s output probabilities for the labels Entailment, Neutral, Contradiction.
+The probabilities are mapped to a final score for each unit as follows:
+[…]
+Finally, we aggregate the individual citation scores to produce three final metrics for the survey: Document-level Accuracy ($R_{d}$), Section-level Accuracy ($R_{s}$), and Sentence-level Accuracy ($R_{t}$).
+For each level $x\in\{d,s,t\}$, the score $R_{x}$ is calculated as the mean of the individual citation scores $y_{x}(r)$ over all citation instances $\mathcal{R}$ in the survey:
+$R_{x}=\frac{1}{|{R}|}\sum_{r\in{R}}y_{x}(r),\quad x\in\{d,s,t\},$ (3)
 4.3 Structural Quality
-A clear and logical structure is fundamental to a high-quality scientific survey, as it greatly enhances readability, coherence, and the ability to quickly navigate information.
-In contrast, poor structural organization can cause confusion, redundancy, and difficulty in following the paper’s arguments.
-Therefore, a robust evaluation of structural quality is essential.
-We propose to assess structure from two complementary perspectives: coarse-grained, high-level organization and fine-grained, heading-level alignment.
-To this end, we introduce two metrics for structural assessment:
-  • Structure Quality Score (SQS): Assesses the high-level structural alignment by comparing the list of generated headings with the ground-truth (GT) headings, focusing on their similarity in structure, meaning, and wording.
-  • Soft-Heading Recall (SHR): Measures the fine-grained alignment of headings and subheadings by using semantic embeddings to account for partial or paraphrased matches.
+The logical flow and coherence of a scientific survey are fundamentally determined by its structure.
+This makes structural quality a critical factor for readability and overall impact.
+To comprehensively evaluate structural quality, we introduce two complementary metrics that assess the generated outline at both macroscopic and microscopic levels.
+Our first metric, the Structure Quality Score (SQS), addresses the high-level organization.
+It holistically assesses the alignment between the generated and ground-truth outlines by comparing their overall structure, semantic coherence, and topical progression.
+Complementing this, our second metric, Soft-Heading Recall (SHR), provides a fine-grained evaluation of heading alignment.
+It specifically measures how well the generated headings cover those in the ground-truth based on the similarity of semantic embeddings.
 Structure Quality Score (SQS).
-SQS is designed to evaluate the overall quality of the generated survey’s outline.
-For this evaluation, we define an outline as a hierarchical list of section and subsection headings extracted from a survey.
-SQS assesses the high-level structural alignment by comparing the generated outline against the ground-truth (GT) outline, focusing on aspects like topical coverage, logical flow, and hierarchical organization.
-To compute this score, we employ the "LLM-as-a-Judge" paradigm.
-[^3] The corresponding prompt templates are available in our official GitHub repository: https://github.com/oneal2000/SurGE.
-Specifically, we prompt a powerful large language model, GPT-4o, with both the generated and the GT outlines.
-The model is then instructed to provide a holistic quality score from 1 (poor) to 5 (excellent).
+SQS evaluates the overall quality of a generated survey’s structure based on the hierarchical list of its section headings.
+To compute this score, we adopt the LLM-as-a-Judge paradigm, leveraging GPT-4o as the evaluator.
+Specifically, we provide the LLM with both the generated and ground-truth outlines and prompt it to assign a quality score.
+To guide the LLM’s evaluation, we have carefully designed a detailed instruction prompt that includes a comprehensive scoring rubric on a scale from 0 to 5.
+The complete prompt is shown in Appendix E.
 Soft-Heading Recall (SHR).
@@
 Formally, SHR is defined as the soft cardinality overlap between the predicted heading set ($H_{P}$) and the ground-truth heading set ($H_{GT}$):
-SHR $=\frac{\mathcal{S}(H_{P}\cap H_{GT})}{\mathcal{S}(H_{GT})}$ (3)
-where $\mathcal{S}(A)$ denotes the soft cardinality of a heading set $A$. It is designed to down-weight redundant headings within the same set and is calculated as:
-$\mathcal{S}(A)$ $=\sum_{i=1}^{K}\frac{1}{\sum_{j=1}^{K}\text{sim}(A_{i},A_{j})}$ (4)
-Here, $\text{sim}(A_{i},A_{j})$ is the cosine similarity between the embeddings of headings $A_{i}$ and $A_{j}$:
-$\text{sim}(A_{i},A_{j})$ $=\cos(\text{embed}(A_{i}),\text{embed}(A_{j}))$ (5)
-Since a standard set intersection would be too strict for comparing paraphrased headings, we calculate the soft cardinality of the intersection $\mathcal{S}(H_{P}\cap H_{GT})$ using the soft set inclusion-exclusion principle:
-$\mathcal{S}(H_{P}\cap H_{GT})=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})-\mathcal{S}(H_{P}\cup H_{GT})$ (6)
-This formulation ensures that semantically matched headings contribute positively to the recall, even with lexical variations.
-A higher SHR score indicates better structural alignment with the GT outline at a granular level.
+SHR $=\frac{\mathcal{S}(H_{P}\cap H_{GT})}{\mathcal{S}(H_{GT})},$ (4)
+where $\mathcal{S}(A)$ denotes the “soft cardinality” of a heading set $A$.
+Intuitively, this metric counts the number of semantically unique headings in a set.
+It achieves this by down-weighting redundant headings.
+Specifically, the contribution of each heading is inversely proportional to its aggregated similarity with all other headings in the set:
+$\mathcal{S}(A)$ $=\sum_{i=1}^{K}\frac{1}{\sum_{j=1}^{K}\text{sim}(A_{i},A_{j})}.$ (5)
+Here, $\text{sim}(A_{i},A_{j})$ is the cosine similarity between the embeddings of headings $A_{i}$ and $A_{j}$.
+A standard set intersection would be too strict for comparing paraphrased headings.
+Therefore, we define the soft intersection cardinality using the inclusion-exclusion principle:
+$\mathcal{S}(H_{P}\cap H_{GT})=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})-\mathcal{S}(H_{P}\cup H_{GT}).$ (6)
+The core idea lies in the union term, $\mathcal{S}(H_{P}\cup H_{GT}$.
+When computed on the combined heading set, a predicted heading and a similar ground-truth heading mutually reduce the union’s soft cardinality.
+This reduction directly quantifies their semantic overlap, allowing the metric to reward paraphrased matches.
+A higher SHR score thus indicates better granular alignment.
 4.4 Content Quality
-Content quality evaluation is pivotal for ensuring surveys are not only comprehensive but also clear, coherent, and logically sound.
-Evaluating this dimension involves assessing the presence of essential information, accuracy, and fluency of expression.
-While traditional metrics such as ROUGE and BLEU effectively measure textual overlap and fidelity, they inadequately capture nuanced aspects of logical coherence and clarity.
-To address these limitations, we supplement these traditional metrics with a comprehensive logical coherence evaluation.
-Metric Definitions.
-We apply ROUGE and BLEU at the section-level to effectively capture fidelity and precision in content representation without needing explicit mathematical formulations here.
-Logical coherence is evaluated using GPT-4o to assign readability and coherence scores (0–5) for each section.
-[^4] Here, we employ the LLM-as-a-Judge paradigm for evaluation.
-The corresponding prompt templates are available on our official GitHub repository: https://github.com/oneal2000/SurGE.
-5 Baseline Implementation
-In this section, we detail the implementation of baseline methods.
-Each baseline follows a two-stage pipeline: (1) retrieving a set of potentially relevant papers for a given topic, and (2) organizing and summarizing the retrieved content to produce a structured survey.
-To ensure a fair comparison, all baselines rely on a shared dense retriever for the first stage.
-In the following subsections, we first describe the training procedure of the shared Paper Retriever, followed by the specific methodologies adopted by each baseline in generating the final survey.
+To assess the content quality of generated scientific surveys, we propose the Content Quality Score (CQS) metric based on the LLM-as-a-Judge paradigm, leveraging GPT-4o to evaluate each section of the survey.
+The evaluation is based on five criteria: fluency and coherence, logical clarity, avoidance of redundancy, clarity of description, and absence of errors.
+To guide the LLM’s evaluation, we designed a detailed instruction prompt for the LLM, which is provided in the Appendix E.
+Each section is scored on a scale of 0 to 5, where a higher score reflects superior fluency, logical progression, and clarity.
+Considering the context length limitations of the LLM, we have it score each survey section by section, and the final score is the average of the scores from all sections.
+As supplementary measures, we also compute ROUGE and BLEU scores, which quantify n-gram overlap between the generated and ground-truth surveys.
+While these are well-established metrics in text generation, their role is secondary in our framework, serving as additional checks for content fidelity rather than a primary assessment method.
+5 Experimental Setup
+In this section, we detail the implementation of our experiment.
+Each baseline follows a two-stage pipeline: (1) retrieving a set of potentially relevant papers for a given topic, and (2) organizing and summarizing the retrieved papers to produce a structured survey.
+For fair comparison, all baselines share the same dense retriever for the first stage.
+In the following subsections, we first describe the training of the shared Paper Retriever, followed by the baseline selection and implementation details.
 5.1 Paper Retriever Training
-We employ a dual-encoder architecture for document retrieval, initialized with roberta-base.
-For a query $q$ and a paper abstract $d$, we first construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$.
+We employ a dual-encoder architecture for retrieval, initialized with roberta-base.
+The training process leverages the benchmark dataset, where each topic description $t$ (introduced in §2) serves as the query $q$.
+For a query $q$ and a paper abstract $d$, we construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$.
 Formally, let $X(q)=[\mathrm{CLS}]\,q\,[\mathrm{SEP}]$ and $X(d)=[\mathrm{CLS}]\,d\,[\mathrm{SEP}]$.
-We then feed these tokens into the transformer encoder to obtain a hidden vector at the $\mathrm{CLS}]$ position:
+We then feed these tokens into roberta-base to obtain the contextualized embedding of the $\mathrm{CLS}]$ token:
 $\mathrm{Emb}(X)=\mathrm{transformer}_{[\mathrm{CLS}]}(X).$ (7)
-The similarity score between the query and the document is computed as the dot product of their embeddings:
+The similarity score between the $q$ and the $d$ is computed as the dot product of their embeddings:
 $S(q,d)=\mathrm{Emb}(X(q))^{\top}\cdot\mathrm{Emb}(X(d)).$ (8)
-During training, each query $Q$ is paired with a relevant document $d^{+}$ to form a positive sample, while negative samples $d^{-}\in N$ are drawn from the corpus.
+During training, each query $Q$ is paired with the relevant documents $d^{+}$ from the ground truth paper to form the positive samples, while negative samples $d^{-}\in N$ are randomly sampled from the corpus.
 The retriever is optimized via the softmax cross-entropy loss:
@@
 This objective encourages the model to assign higher scores to relevant papers while minimizing scores for irrelevant ones.
-We use the trained retriever to retrieve the top-ranked papers for each query, thereby providing a focused subset of literature for subsequent summarization.
-5.2 Retrieval-Augmented Generation
-In the first baseline, we combine retrieval with a direct generation approach.
-Given a topic $t$, we use the above retriever to collect the top 100 candidate papers.
-To manage lengthy inputs, these retrieved papers are split into smaller groups, each containing an approximately equal number of references.
-We then prompt a large language model (LLM) to summarize each group separately, guiding it to preserve references to the original papers.
-Formally, for each group of papers $G_{k}=\{d_{1},d_{2},\ldots,d_{n_{k}}\}$, the LLM is conditioned on the sequence of paper abstracts and instructed to produce a partial summary $\hat{S}_{k}$.
-Finally, we merge the partial summaries $\hat{S}_{1},\hat{S}_{2},\ldots$ into a unified survey.
-This merging step is performed by prompting the LLM once again with all partial summaries and asking for an integrated, logically coherent survey.
-Although this baseline follows a straightforward two-step approach, it provides a clear assessment of how effective retrieval-based summarization can be when coupled with an LLM’s generative capabilities.
-5.3 AutoSurvey
-AutoSurvey wang2024autosurvey implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions.
-We adapt it to use our fine-tuned retriever in place of its original retrieval mechanism and keep the number of retrieved references consistent for fairness.
-In the adapted workflow, we first issue a query based on the topic $t$ to retrieve an initial collection of papers $P_{\text{init}}$.
-The LLM is then prompted to create a structured outline, which includes main sections and subsections tailored to the subject matter.
-Next, each section is expanded by conditioning on the subset of papers most relevant to that specific section, producing a draft that includes references in a bracketed format (e.g., “[id]”).
-Once each section is drafted, the LLM refines it to address factual inconsistencies, stylistic mismatches, and reference-formatting issues.
-Finally, all refined sections are merged into a coherent final survey, with transitions and citation references carefully aligned.
-The workflow iterates over these stages, leading to incremental improvements in both thematic coverage and presentation quality.
-5.4 StepSurvey
-StepSurvey lai2024instruct is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner.
-This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 tian2024overview .
-Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature.
-Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers.
-The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections.
-Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy.
-By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
+After training, we use the trained retriever to retrieve the top-ranked papers for each query, thereby providing a collection of relevant papers for the subsequent generation stage.
+5.2 Baselines
+We selected the following three survey generation baseline methods for our experiments. Detailed descriptions of each method are provided in the appendices.
+  • Retrieval-Augmented Generation (RAG): This straightforward approach retrieves relevant papers, summarizes them in chunks, and then merges these summaries to form the final survey.
+A detailed description is provided in Appendix D.1.
+  • AutoSurvey (Wang et al., 2024d): This method employs a multi-stage, outline-driven pipeline.
+It first generates a high-level outline from retrieved papers and then iteratively expands and refines each section.
+Further details can be found in Appendix D.2.
+  • StepSurvey (Lai et al., 2024b): This approach adopts a granular, step-by-step process.
+It starts by generating a title and primary headings, which then guide the incremental drafting of each subsection.
+The methodology is described in detail in Appendix D.3.
 5.3 Implementation Details.
 For the training of Paper Retriever, we randomly split the dataset into a training set and a test set at a ratio of 4:1.
-We adopt the AdamW optimizer for model optimization, the learning rate is set to $5\times 10^{-6}$, the weight decay is set to $1\times 10^{-2}$, and the epoch is set to $10$.
-During the training process, we adopt mixed-precision (fp16) training for enhanced efficiency.
-At inference time, each query retrieves the top 100 most relevant papers according to the similarity score $S(q,d)$.
-The retriever is initialized using the pre-trained RoBERTa model liu2019roberta .
+We adopt the AdamW optimizer for model optimization, the learning rate is set to $5\times 10^{-6}$, and the epoch is set to $10$.
+During the training process, we adopt mixed-precision (fp16) training.
+At inference time, each query retrieves the top 100 relevant papers according to the similarity score.
+The retriever is initialized using the pre-trained RoBERTa model (Liu et al., 2019).
+To ensure a fair comparison, we utilize the Qwen2.5-14B-Instruct model (Yang et al., 2024a) as the base LLM for all baseline methods.
 For the generation configuration of LLMs, all experiments are conducted using the publicly available implementations provided by Hugging Face.
 We utilize the default hyperparameters and the chat template as outlined in the official Hugging Face repository  [^4] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
-All experiments are conducted on a server with 1TB of RAM and eight NVIDIA A100 GPUs, each with 40GB of memory.
+All experiments are conducted on a GPU server with 1TB RAM and eight NVIDIA A100 GPUs, each with 40GB of memory.
 6 Experimental Results
-This section demonstrates the utility of our proposed benchmark by presenting a detailed evaluation of three representative survey generation baselines: RAG, AutoSurvey, and StepSurvey.
-Our analysis is designed to showcase how the benchmark’s multi-faceted metrics can be used to conduct a fine-grained analysis and pinpoint system bottlenecks.
-We first isolate and evaluate the performance of the shared Paper Retriever (§6.1).
-This analysis is crucial as it establishes the theoretical upper-bound performance for the reference coverage metric and highlights the challenges posed by the initial retrieval stage.
-Following this, we assess the complete end-to-end performance of the baselines (§6.2), leveraging our comprehensive evaluation suite to compare their architectural strengths and weaknesses and to shed light on the broader challenges facing automatic survey generation.
+This section presents the detailed evaluation of the selected survey generation baselines.
+We first evaluate the performance of the shared Paper Retriever (§6.1).
+This analysis is crucial as it establishes the theoretical upper-bound performance for the comprehensiveness metric and highlights the challenges posed by the initial retrieval stage.
+Following this, we assess the complete end-to-end performance of the three baselines (§6.2) based on our proposed evaluation framework.
 Table 1: Comparison of retrieval models on recalling ground-truth cited papers. The metric is Recall@k, where $k$ is the number of top documents retrieved. Best results are in bold.
@@
 A crucial question in our two-stage pipeline is whether performance limitations stem from the retriever’s inability to find relevant papers or the generator’s inability to use them.
-To disentangle these factors and quantify the retrieval bottleneck, we first evaluate the performance of our fine-tuned dense retriever in isolation.
+To disentangle these factors and quantify the retrieval bottleneck, we evaluate the performance of our fine-tuned dense retriever in isolation.
 This analysis establishes the theoretical upper bound for the reference coverage that our end-to-end systems can achieve.
-We compare our dense retriever against a lexical baseline BM25 robertson2009probabilistic , using Recall@k as the evaluation metric.
+We compare our dense retriever against the lexical baseline BM25 (Robertson et al., 2009), using Recall@k as the evaluation metric.
 This metric measures the percentage of ground-truth papers from the reference survey that are present in the top-k retrieved documents.
@@
 The performance gap underscores the inadequacy of lexical search, which struggles to find semantically relevant papers that may not share overlapping keywords.
-Our dense retriever, by capturing deeper semantic relationships, is far more effective.
+On the other hand, our dense retriever can capture deeper semantic relationships, thus it is far more effective than BM25.
 However, the results also reveal a critical bottleneck in the survey generation pipeline.
-Even with a generous budget of k=1000, Paper Retriever recalls only 68.05% of the ground-truth papers.
-This finding is crucial as it establishes a hard upper-bound on the performance of the downstream generator.
-This finding suggests that achieving a breakthrough in survey quality may necessitate a shift towards more sophisticated retrieval paradigms, such as employing search agents powered by large language models zhang2024large .
-Table 2: The experimental results combining Coverage, Referencing Accuracy (including Doc-Rel, Sec-Rel, and Sent-Rel), Structural Quality (including SQS and SHR), and Content Quality (as measured by ROUGE scores, BLEU, and Logic Score).
-Doc-Rel represents document-level relevance, Sec-Rel represents section-level relevance, Sent-Rel represents sentence-level relevance, SQS represents the structure quality score, and SHR represents soft-heading recall.
-R-L represents ROUGE-L.
+Even when retrieving the top 1000 documents (k=1000), Paper Retriever’s recall reaches only 68.05% of the ground-truth papers.
+This performance ceiling imposes a hard upper bound on the downstream generator, as it cannot synthesize information from papers it never receives.
+Consequently, this performance ceiling underscores that the survey generation task needs more sophisticated retrieval paradigms, such as employing search agents or multi-agent systems powered by large language models (Zhang et al., 2024; Tang et al., 2025; Li et al., 2025).
+Table 2: Main experimental results comparing different survey generation baselines across four dimensions: Comprehensiveness (Comp.), Citation Accuracy, Structural Quality, and Content Quality.
+Metrics include Recall, Document/Section/Sentence-level Citation Accuracy (Doc-Acc, Sec-Acc, Sent-Acc), Structure Quality Score (SQS), Soft-Heading Recall (SHR), ROUGE-L, BLEU, and Content Quality Score (CQS).
 The best results are in bold, and the second-best results are underlined.
-Coverage Referencing Accuracy Structural Quality Content Quality
-Baseline Recall Doc-Rel Sec-Rel Sent-Rel SQS SHR R-L BLEU Logic Score
+Comp. Citation Accuracy Structural Quality Content Quality
+Baseline Recall Doc-Acc Sec-Acc Sent-Acc SQS SHR R-L BLEU CQS
 RAG 0.0214 0.2857 0.2502 0.2500 0.6829 0.7900 0.1519 10.38 4.6723
@@
 6.2 Overall Performance of Survey Generation Baselines
-The preceding analysis revealed that the retriever provides the generation models with over a third (36.65%) of the ground-truth references, setting a clear performance ceiling.
-We now demonstrate the full diagnostic power of our benchmark by applying its comprehensive suite of metrics to the end-to-end outputs of the three baselines: RAG, AutoSurvey, and StepSurvey.
-As detailed in Table 2, this multi-faceted evaluation allows us to pinpoint specific challenges in the generation stage.
-The key insights revealed by our benchmark are as follows:
-(1) Bottleneck of Existing Baselines.
-A striking observation across all methods is the significant gap between retrieval performance and final survey coverage.
-While our paper retriever makes 36.65% of the ground-truth references available (Recall@100), the best-performing baseline, StepSurvey, only achieves a final Coverage of 6.30%.
-The simpler RAG baseline is even lower at 2.14%.
-This disparity underscores that the primary limitation is the generation stage’s inability to effectively identify and incorporate relevant information from the provided set of 100 papers.
-Nonetheless, the results also show that advanced, structured approaches like AutoSurvey (3.51%) and StepSurvey (6.30%) are substantially better at utilizing the retrieved context than a standard RAG pipeline.
-(2) Strengths of AutoSurvey.
-AutoSurvey exhibits excellent Section-Level Relevance (0.4935) and Sentence-Level Relevance (0.4870), surpassing both RAG and StepSurvey.
-This improvement appears to stem from AutoSurvey’s iterative planning mechanism, where the system generates a high-level outline, retrieves relevant references for each section, and incrementally refines each subsection.
-By focusing on context-specific references and refining drafts section by section, AutoSurvey places citations more accurately at both the section and sentence levels.
-Moreover, AutoSurvey achieves the highest Structure Quality Score (SQS) of 1.3902, indicating that its top-down planning leads to well-organized hierarchical structures.
-(3) Strengths of StepSurvey.
-On the other hand, StepSurvey attains the best performance in several crucial dimensions: Coverage (6.30%) and Document-Level Relevance (0.4576).
-Its multi-phase workflow first generates major headings, then subtopics, and finally assigns references to each subsection.
-This step-by-step refinement appears particularly effective at broadening the coverage of included works and aligning the cited literature with the overarching survey theme.
-StepSurvey also achieves the highest Logic Score (4.8451) and excels in ROUGE and BLEU, reflecting strong content quality and coherent presentation.
-An interesting observation is that while StepSurvey’s Section-Level Relevance is slightly lower than AutoSurvey’s, the difference remains small (0.4571 vs.
-0.4935), indicating that its primary strength lies in more holistic, document-wide alignment and consistent logical flow.
-(4) Impact of Planning on Multi-Level Relevance.
-The multi-stage generation process adopted by AutoSurvey and StepSurvey benefits Document-Level and Section-Level Relevance in different ways.
-AutoSurvey’s approach—outlining, drafting, and iteratively refining sections—yields highly relevant citations at the subsection and sentence granularity.
-In contrast, StepSurvey’s method of organizing subtopics around central themes ensures strong document-wide alignment.
-This discrepancy suggests that carefully balancing local vs.
-global planning remains an open challenge: future work might explore hybrid strategies that combine AutoSurvey’s fine-grained focus with StepSurvey’s broad thematic coverage.
-In conclusion, our results demonstrate that advanced, multi-stage planning approaches (AutoSurvey and StepSurvey) significantly outperform the standard RAG pipeline in nearly all metrics.
-While AutoSurvey excels in section- and sentence-level citation placement and high-level structural quality, StepSurvey provides superior document-level coverage, global relevance, and strong logical coherence.
-Nonetheless, all methods still struggle to capture the vast reference space fully.
-These findings underscore the importance of iterative refinement and structured planning in automated survey generation and suggest that there remains substantial room for further improvement, particularly in boosting coverage and refining the interplay between local and global survey organizations.
+Our evaluation now transitions from retrieval performance to the end-to-end survey generation task.
+While the retriever supplies the models with a substantial portion (36.65%) of the ground-truth references, a detailed analysis of the final surveys reveals a significant bottleneck in the generation stage.
+Table 2 presents a multi-faceted comparison of our three baselines across the four dimensions of survey quality defined by our benchmark.
+This comprehensive evaluation yields several key insights:
+(1) The Generation Stage is the Primary Bottleneck.
+A striking finding is the dramatic performance drop-off between paper retrieval and final survey synthesis.
+Despite the retriever providing access to 36.65% of ground-truth references (Recall@100), the best-performing model, StepSurvey, incorporates only 6.30% of them into its final output.
+The standard RAG baseline fares even worse, with a final recall of just 2.14%.
+This stark disparity highlights that the core challenge lies not in finding relevant papers, but in the generator’s ability to effectively identify and integrate crucial information from the retrieved set.
+Nevertheless, the results also affirm the value of structured pipelines, as AutoSurvey (3.51%) and StepSurvey (6.30%) significantly outperform the naive RAG approach in comprehensiveness.
+(2) AutoSurvey Excels in Local Coherence and Structural Quality.
+AutoSurvey demonstrates particular strength in producing locally coherent and well-organized content.
+As shown in Table 2, it achieves the highest scores for both Section-level Citation Accuracy (Sec-Acc: 0.4935) and Sentence-level Citation Accuracy (Sent-Acc: 0.4870).
+This suggests its iterative, section-by-section refinement process is highly effective at placing citations within their correct immediate context.
+Furthermore, their planning approach results in the highest Structure Quality Score (SQS) of 1.3902, indicating superior hierarchical organization in the generated survey.
+(3) StepSurvey Achieves Superior Coverage and Global Content Quality.
+In contrast, StepSurvey’s strengths lie in metrics related to holistic coverage and overall content quality.
+It attains the highest Recall (0.0630) and Document-level Citation Accuracy (Doc-Acc: 0.4576), demonstrating a better ability to cover the breadth of the topic and align cited works with the survey’s main theme.
+Its multi-phase workflow that generates headings, then subtopics, and finally assigns citations leads to the best content quality.
+This is evidenced by its top scores in text similarity metrics (ROUGE-L: 0.1590, BLEU: 12.02) as well as the highest Content Quality Score (CQS) of 4.8451, which reflects stronger logical flow and presentation.
+(4) A Trade-off Between Local Precision and Global Coverage.
+The distinct strengths of AutoSurvey and StepSurvey reveal a fundamental trade-off in planning strategies.
+AutoSurvey’s iterative refinement of individual sections fosters high-quality local structure and precise, fine-grained citation placement.
+Conversely, StepSurvey’s hierarchical, topic-first approach yields better overall topic coverage and more coherent, globally relevant content.
+While both advanced methods significantly outperform the standard RAG pipeline, neither excels across all dimensions.
+In conclusion, our proposed SurGE benchmark effectively diagnoses the weaknesses of current survey generation systems.
+The results clearly indicate that advanced, multi-stage planning is crucial, but a substantial gap remains between machine-generated and expert-written surveys, particularly in comprehensiveness.
+The observed trade-off between local and global optimization strategies highlights a key challenge for future research: developing hybrid models that can combine the fine-grained accuracy of AutoSurvey with the broad thematic coverage and coherence of StepSurvey.
 7 Related Work
@@
 Large Language Models (LLMs) are inherently limited by their static, pre-trained parametric knowledge.
-To address these limitations, Retrieval-Augmented Generation (RAG) has emerged as a key paradigm gao2023retrieval ; lewis2020retrieval ; dong2025decoupling ; tu2025rbft ; su2025dynamic .
-By grounding the model in external knowledge, RAG directly addresses several fundamental limitations of LLMs, offering a robust mechanism to mitigate hallucinations ji2023survey ; su2024unsupervised ; su2024mitigating ; wang2025joint , facilitate knowledge updating fang2024alphaedit ; wang2024lekube ; wang2024knowledge ; wang2025decoupling , and enable effective domain adaptation yang2024tcm ; su2025judge ; su2024stard ; su2024legalaid .
-The conventional approach to traditional RAG is built upon the "Retrieval-then-Read" paradigm borgeaud2022improving ; guu2020retrieval ; lewis2020retrieval .
+To address these limitations, Retrieval-Augmented Generation (RAG) has emerged as a key paradigm (Gao et al., 2023; Lewis et al., 2020; Dong et al., 2025; Tu et al., 2025; Su et al., 2025a).
+By grounding the model in external knowledge, RAG directly addresses several fundamental limitations of LLMs, offering a robust mechanism to mitigate hallucinations (Ji et al., 2023; Su et al., 2024d; b; Wang et al., 2025a), facilitate knowledge updating (Fang et al., 2024a; Wang et al., 2024b; a; 2025b), and enable effective domain adaptation (Yang et al., 2024b; Su et al., 2025c; 2024a; 2024e).
+The conventional approach to traditional RAG is built upon the ”Retrieval-then-Read” paradigm (Borgeaud et al., 2022; Guu et al., 2020; Lewis et al., 2020).
 Within this framework, a user’s query triggers a search module for relevant documents within a large-scale external corpus.
-This retrieval step is carried out by either an external retriever zhai2008statistical ; su2023caseformer ; robertson2009probabilistic ; su2023wikiformer ; ma2023caseencoder ; fang2024scaling or a more sophisticated retrieval system su2023thuir2 ; salemi2024towards .
+This retrieval step is carried out by either an external retriever (Zhai, 2008; Su et al., 2023b; Robertson et al., 2009; Su et al., 2023a; Ma et al., 2023; Fang et al., 2024b) or a more sophisticated retrieval system (Su et al., 2023c; Salemi & Zamani, 2024; Chen et al., 2022).
 Building upon this foundation, recent work has proposed more advanced RAG architectures to improve efficiency and effectiveness.
-For instance, Dynamic RAG jiang2022retrieval ; su2024dragin ; yao2024seakr moves beyond a single retrieval step by adaptively triggering the retriever during generation, specifically when the LLM is uncertain during the generation process.
-From another angle, GraphRAG edge2024local enhances the knowledge source by querying pre-constructed knowledge graphs instead of unstructured text, allowing it to retrieve interconnected facts and relationships.
-Furthermore, the Parametric RAG paradigm  su2025parametric ; tan2025dynamic ; fleshman2025lora alters the knowledge injection step by directly injecting retrieved knowledge into the LLM’s parameters.
+For instance, Dynamic RAG (Jiang et al., 2022; Su et al., 2024c; Yao et al., 2024) moves beyond a single retrieval step by adaptively triggering the retriever during generation, specifically when the LLM is uncertain during the generation process.
+From another angle, GraphRAG (Edge et al., 2024) enhances the knowledge source by querying pre-constructed knowledge graphs instead of unstructured text, allowing it to retrieve interconnected facts and relationships.
+Furthermore, the Parametric RAG paradigm  (Su et al., 2025b; Tan et al., 2025; Fleshman & Van Durme, 2025) alters the knowledge injection step by directly injecting retrieved knowledge into the LLM’s parameters.
 The scientific survey generation task, which is the focus of our SurGE benchmark, presents a significant challenge for even these advanced RAG systems.
@@
 Long-form text generation is substantially more challenging than short-text generation due to its inherent requirements for sustained coherence and rich contextual understanding.
-Early approaches mainly used generative adversarial networks and reinforcement learning to conduct long-sequence generation (Guo2018, ).
+Early approaches mainly used generative adversarial networks and reinforcement learning to conduct long-sequence generation (Guo et al., 2018).
 More recently, large language models have emerged as a strong tool for this task, offering advanced capabilities to handle long-text generation.
-For example, structured planning techniques and specialized inference mechanisms are proposed (Sloan2025, ; Jin2024, ) to generate consistent and high-quality clinical reports.
-Similarly, hierarchical planning frameworks have demonstrated that content control and multi-constraint instruction following can significantly enhance logical flow and topic coverage (hu2022, ; pham2024, ).
-Beyond medical or other task-specific applications, context-driven retrieval strategies, such as tree-structured retrieval, can support open-domain long-text generation by guiding the model through extensive knowledge sources (roy2024, ).
+For example, structured planning techniques and specialized inference mechanisms are proposed (Sloan et al., 2024; Jin et al., 2024) to generate consistent and high-quality clinical reports.
+Similarly, hierarchical planning frameworks have demonstrated that content control and multi-constraint instruction following can significantly enhance logical flow and topic coverage (Hu et al., 2022; Pham et al., 2024).
+Beyond medical or other task-specific applications, context-driven retrieval strategies, such as tree-structured retrieval, can support open-domain long-text generation by guiding the model through extensive knowledge sources (Roy et al., 2024).
 The effective evaluation framework is vital for measuring the quality, factualness, and user-centric utility of the long-form text generation task.
 Traditional metrics, designed for shorter texts, often fail to capture the intricacies of longer outputs.
-Recent work has introduced task-focused benchmarks that emphasize user-oriented objectives, such as personalized writing or domain-specific content generation (kumar2024, ; salemi2025, ).
+Recent work has introduced task-focused benchmarks that emphasize user-oriented objectives, such as personalized writing or domain-specific content generation (Kumar et al., 2024; Salemi et al., 2025).
 In parallel, factuality assessment has attracted growing interest, with methods proposed to evaluate both verifiable and unverifiable claims.
-Metrics such as VERISCORE and FACTSCORE break down generated text into atomic facts, checking each for consistency against reliable sources (song2024, ; min2023, ).
+Metrics such as VERISCORE and FACTSCORE break down generated text into atomic facts, checking each for consistency against reliable sources (Song et al., 2024; Min et al., 2023).
 Beyond factual correctness, coherence and structural quality have been studied extensively.
-Benchmarks like LongGenBench and HelloBench underscore the importance of evaluating a model’s ability to maintain logical organization and clarity over extended passages (wu2025, ; que2024, ).
+Benchmarks like LongGenBench and HelloBench underscore the importance of evaluating a model’s ability to maintain logical organization and clarity over extended passages (Wu et al., 2025; Que et al., 2024).
 7.3 Survey Generation
 In the domain of scientific writing, survey generation involves distilling extensive textual resources into a coherent and structured overview.
-Recent advances in AI-assisted systems have provided prompting-based approaches to expedite the drafting process while preserving content accuracy (gero2022, ; Kacena2024, ).
-One of the most commonly used approaches is retrieval-augmented generation, which combines large-scale knowledge retrieved from documents with language generation empowered by LLMs to yield factually comprehensive overviews (lewis2021, ).
-Retrieval-augmented generation is often initiated with dense retrieval methods based on dual-encoder architectures to identify highly relevant documents (karpukhin2020, ).
-Once these documents are retrieved, summarization techniques—spanning top-down, bottom-up, and graph-based ranking methods—play a pivotal role in producing concise yet faithful summaries (nayeem2024, ; pang2022, ; bleiweiss2023, ).
-Building on these retrieval and summarization-based methodologies, automated literature survey generation has garnered increasing attention (nlpcc2024, ; lai2024, ).
-However, existing techniques depend on limited ground truths and employ coarse evaluation metrics, resulting in oversimplified assessments of survey quality wang2024autosurvey .
+Recent advances in AI-assisted systems have provided prompting-based approaches to expedite the drafting process while preserving content accuracy (Gero et al., 2022; Kacena et al., 2024).
+One of the most commonly used approaches is retrieval-augmented generation, which combines large-scale knowledge retrieved from documents with language generation empowered by LLMs to yield factually comprehensive overviews (Lewis et al., 2021).
+Retrieval-augmented generation is often initiated with dense retrieval methods based on dual-encoder architectures to identify highly relevant documents (Karpukhin et al., 2020).
+Once these documents are retrieved, summarization techniques—spanning top-down, bottom-up, and graph-based ranking methods—play a pivotal role in producing concise yet faithful summaries (Nayeem & Rafiei, 2024; Pang et al., 2022; Bleiweiss, 2023).
+Building on these retrieval and summarization-based methodologies, automated literature survey generation has garnered increasing attention (Tian et al., 2024a; Lai et al., 2024a).
+However, existing techniques depend on limited ground truths and employ coarse evaluation metrics, resulting in oversimplified assessments of survey quality (Wang et al., 2024d).
 To address these challenges, we present a refined ground truth and a multi-dimensional evaluation framework that emphasizes both accuracy and structural coherence.
@@
 8 Conclusion
-In this paper, we introduce SurGE, a benchmark for end-to-end scientific survey generation in computer science.
-We construct a high-quality dataset and develop a multi-dimensional evaluation framework that assesses the coverage, coherence, factual accuracy, and writing quality of automatically generated surveys.
-Extensive experiments with various baselines highlight the challenges in producing high-quality surveys.
-SurGE is expected to serve as a valuable resource for future research in information retrieval, fostering advancements in both retrieval methodologies and generative AI techniques.
-Several limitations guide future work, including: (1) The benchmark is limited in the domain of computer science but could be extended to broader domains.
-(2)  Integration with more advanced retrieval and generation methods could further enhance automated scientific survey generation for computer science.
-Appendix A Details of the Academic Corpus
+In this paper, we introduce SurGE, a comprehensive benchmark designed to address the critical need for standardized and reproducible evaluation in automated scientific survey generation.
+SurGE provides a large-scale academic corpus, a set of expert-written ground-truth surveys, and a fully automated framework to evaluate surveys on comprehensiveness, citation accuracy, structure, and content.
+Our experiments reveal significant limitations in state-of-the-art LLM-based systems, highlighting challenges such as incomplete topic coverage and reference hallucination.
+We believe SurGE will catalyze future research and guide the development of more effective LLM-based systems at the intersection of information retrieval and generative AI for this important task.
+9 Ethics Statement
+The SurGE benchmark aims to facilitate the advancement of automated survey generation and evaluation in the context of scientific research.
+By utilizing publicly available metadata from arXiv, our work prioritizes transparency and open access.
+We do not host or redistribute arXiv’s PDF content, ensuring compliance with arXiv’s Terms of Use and the CC0 public-domain dedication for metadata.
+While the use of large-scale academic data in SurGE poses minimal ethical risks, we emphasize the importance of ethical data usage, privacy, and responsible development of AI systems.
+We encourage the research community to consider these aspects when applying or building upon SurGE.
+10 Reproducibility Statement
+To ensure reproducibility, all code, data, and models for the SurGE benchmark are open-sourced and available at our official GitHub Repository [^5] https://github.com/oneal2000/SurGE.
+Detailed descriptions of the dataset construction, processing steps, and evaluation methods can be found at our official GitHub Repository and appendices.
+Our open-access approach is designed to facilitate the replication of results and encourage future research in the field of survey generation using large language models.
+Appendix A Reference Extraction and Processing
+[…]
+The core reason for extracting the full reference list from each ground-truth survey is to create a gold standard for evaluating the information collection stage.
+We operate on the premise that the citations in an expert-written survey represent a curated collection of the field’s most foundational and relevant literature.
+By using this set as the ground truth, we can quantitatively measure the coverage and recall of a system-generated survey’s references, providing a clear metric for its performance.
+Appendix B Details of the Academic Corpus
 Table 3 provides an overview of the metadata schema used in our academic corpus. Each entry is structured to support efficient retrieval and interpretability.
@@
 doc_id A unique identifier assigned for reference and retrieval.
-Appendix B Social Impact
-Automated scientific survey writing brings two-sided social impacts.
-On the one hand, automated survey generation can save significant time and resources for researchers.
-This efficiency allows for quick data collection and analysis, enabling more rapid responses to a research domain.
-Automated survey generation could even provide innovation and lead to discoveries and insights with the help of LLMs.
-On the other hand, we have shown that the current method for automated survey generation is limited.
-Such a low-quality survey could poison the academic environment and bring ethical concerns, such as biased content generated by LLMs.
-Therefore, surveys generated by LLMs should be noted and published in different tracks.
-Such research direction should be carefully supervised by the collaboration of researchers.
-Appendix C Selected Ground Truth Survey
-The following three tables provide the complete list of the ground-truth surveys that comprise the SurGE benchmark.
-In line with our selection methodology outlined in Section 3, these surveys represent high-impact, peer-reviewed works.
-For each entry, we list its full title, publication year, primary arXiv category, and its citation count.
-The provided citation counts represent a snapshot taken from Google Scholar on May 10, 2025.
+Appendix C The Use of Large Language Models
+Large language models were used as a tool to help refine the language and improve the clarity of this manuscript.
+Following any such use, the authors reviewed and edited the content to ensure its accuracy and originality, and take full responsibility for the final text.
+Appendix D Baseline Implementation
+D.1 Retrieval-Augmented Generation
+In the first baseline, we combine retrieval with a direct generation approach.
+Given a topic $t$, we use the above retriever to collect the top 100 candidate papers.
+To manage lengthy inputs, these retrieved papers are split into smaller groups, each containing an approximately equal number of references.
+We then prompt a large language model (LLM) to summarize each group separately, guiding it to preserve references to the original papers.
+Formally, for each group of papers $G_{k}=\{d_{1},d_{2},\ldots,d_{n_{k}}\}$, the LLM is conditioned on the sequence of paper abstracts and instructed to produce a partial summary $\hat{S}_{k}$.
+Finally, we merge the partial summaries $\hat{S}_{1},\hat{S}_{2},\ldots$ into a unified survey.
+This merging step is performed by prompting the LLM once again with all partial summaries and asking for an integrated, logically coherent survey.
+Although this baseline follows a straightforward two-step approach, it provides a clear assessment of how effective retrieval-based summarization can be when coupled with an LLM’s generative capabilities.
+D.2 AutoSurvey
+AutoSurvey (Wang et al., 2024d) implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions.
+We adapt it to use our fine-tuned retriever in place of its original retrieval mechanism and keep the number of retrieved references consistent for fairness.
+In the adapted workflow, we first issue a query based on the topic $t$ to retrieve an initial collection of papers $P_{\text{init}}$.
+The LLM is then prompted to create a structured outline, which includes main sections and subsections tailored to the subject matter.
+Next, each section is expanded by conditioning on the subset of papers most relevant to that specific section, producing a draft that includes references in a bracketed format (e.g., “[id]”).
+Once each section is drafted, the LLM refines it to address factual inconsistencies, stylistic mismatches, and reference-formatting issues.
+Finally, all refined sections are merged into a coherent final survey, with transitions and citation references carefully aligned.
+The workflow iterates over these stages, leading to incremental improvements in both thematic coverage and presentation quality.
+D.3 StepSurvey
+StepSurvey (Lai et al., 2024b) is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner.
+This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 (Tian et al., 2024b).
+Rather than producing an overarching outline at once, it starts by proposing a survey title and a set of primary headings that collectively capture the central themes of the retrieved literature.
+Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers.
+The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections.
+Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy.
+By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
+Appendix E Prompt Template
+This section provides the detailed prompt templates used for our LLM-as-a-Judge evaluation, as described in the main body of the paper.
+These prompts are specifically designed for GPT-4o to assess the Structure Quality Score (SQS) and the Content Quality of the generated surveys.
+Prompt Template for Structure Quality Score You are an AI evaluator.
+Your task is to compare the generated titles with the target titles and assign a score from 0 to 5 based on their similarity in structure, meaning, and wording.
+Target Titles: {Ground Truth Titles} Generated Titles: {Generated Titles} Scoring Criteria: 5 – Almost Identical: Nearly all key words match exactly.
+The meaning is fully preserved.
+The phrasing and structure are identical or differ only in trivial ways.
+4 – Very Similar: Most key words match.
+The meaning is nearly identical.
+The phrasing and structure are very close, with minor rewording.
+3 – Similar: Several key words are shared.
+The meaning is largely the same with slight variations.
+The structure is somewhat similar, but there may be word substitutions.
+2 – Somewhat Similar: Some key words are shared, but others are different.
+The general topic is the same, but the emphasis may differ.
+The sentence structures are different but not entirely unrelated.
+1 – Somewhat Different: Few words overlap, but they are not key terms.
+The meaning is somewhat related but mostly different.
+The sentence structures are significantly different.
+0 – Completely Different: Nearly no words in common.
+Completely different meanings.
+No similarity in structure or phrasing.
+Instruction: Analyze the generated titles based on the criteria above and provide a single score between 0 and 5.
+Prompt Template for Content Quality Score You are an advanced AI language evaluator.
+Your task is to assess the logical coherence and clarity of the text based on the following criteria: 1.
+Fluency and Coherence – Does the text flow naturally?
+Are the sentences well-connected and easy to read?
+2.
+Logical Clarity – Is the reasoning clear and structured?
+Does the argument progress logically without contradictions?
+3.
+Avoidance of Redundancy – Does the text avoid unnecessary repetition?
+4.
+Clarity of Description – Are ideas, concepts, or events described in a way that is easy to understand?
+5.
+Absence of Errors – Does the text contain grammatical mistakes, spelling errors, or factual inconsistencies?
+You will provide a score from 0 to 5 based on the following criteria: 5 – Excellent: A score of 5 is awarded to texts that are highly fluent, featuring smooth transitions and a natural flow.
+The logical progression is clear, well-structured, and easy to follow.
+There is no redundancy; each sentence contributes meaningfully to the overall message.
+Furthermore, descriptions are precise and unambiguous, and the text is free of any spelling, grammatical, or factual errors.
+4 – Good A score of 4 indicates a text that is mostly fluent but may have minor awkward transitions.
+Its logical progression is clear, though it might contain slight inconsistencies.
+There may be some minor redundancy or repetition present.
+While descriptions are mostly clear, they could contain minor ambiguities, and the text has very few spelling or grammatical errors.
+3 – Average A score of 3 applies to texts that are understandable despite containing noticeable awkward phrasing.
+The logical flow is inconsistent, which may cause some points to feel out of place.
+Some redundancy or repetition is present and slightly affects readability.
+Additionally, certain descriptions are vague or unclear, and the text contains some spelling or grammatical mistakes but remains readable.
+2 – Poor A score of 2 is assigned when a text is difficult to read due to an awkward structure and poor fluency.
+Logical inconsistencies make the central argument unclear, and repetitive phrases render the content tedious.
+Descriptions are vague, making it hard to understand key points, and the submission is characterized by multiple grammatical and spelling errors.
+1 – Very Poor A score of 1 denotes a text that is highly disjointed, making it very hard to read.
+The logical flow is almost nonexistent, with abrupt topic shifts throughout.
+Redundant sentences are included but add no value to the content.
+Descriptions are confusing or overly vague, and the text suffers from frequent spelling and grammatical mistakes.
+0 – Incoherent A score of 0 is reserved for a text that is completely nonsensical or unreadable.
+It demonstrates no logical progression or coherence.
+The content exhibits extreme redundancy or devolves into word salad, and severe errors throughout make it impossible to understand any intended meaning.
+Instruction Now evaluate the following paragraph based on the criteria above and provide a score from 0 to 5.
+Paragraph: {Paragraph}
+Appendix F Selected Ground Truth Survey
+The following three tables provide the complete list of the ground-truth surveys in the SurGE benchmark.
+For each entry, we list the title, publication year, primary arXiv category, and citation count.
+The provided citation counts represent a snapshot from Google Scholar on May 10, 2025.
 Survey Title Year Category Citation Count
@@
 Deep Learning for Vision-based Prediction: A Survey 2020 cs.CV 59
-A Survey of Machine Learning for Computer Architecture and Systems 2021 cs.LG 91
-Inspect, Understand, Overcome:A Survey of Practical Methodsfor AI Safety 2021 cs.LG 78
-Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394
-A Survey on Recent Approaches for Natural Language Processing in Low-Resource Scenarios 2020 cs.CL 390
-DL-Traff: Survey and Benchmark of Deep Learning Models for Urban Traffic Prediction 2021 cs.LG 180
 Survey Title Year Category Citation Count
@@
 A Survey on Dialogue Summarization: Recent Advances and New Frontiers 2021 cs.CL 111
+A Survey of Machine Learning for Computer Architecture and Systems 2021 cs.LG 91
+Inspect, Understand, Overcome: A Survey of Practical Methods for AI Safety 2021 cs.LG 78
+Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394
+A Survey on Recent Approaches for Natural Language Processing in Low-Resource Scenarios 2020 cs.CL 390
+DL-Traff: Survey and Benchmark of Deep Learning Models for Urban Traffic Prediction 2021 cs.LG 180

### v4

@@
 Abstract
-The rapid growth of academic literature makes the manual creation of scientific surveys increasingly infeasible.
-While large language models show promise for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols.
-To bridge this critical gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for scientific survey generation in computer science.
-SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers.
-In addition, we propose an automated evaluation framework that measures the quality of generated surveys across four dimensions: comprehensiveness, citation accuracy, structural organization, and content quality.
-Our evaluation of diverse LLM-based methods demonstrates a significant performance gap, revealing that even advanced agentic frameworks struggle with the complexities of survey generation and highlighting the need for future research in this area.
-[^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
+The exponential growth of scientific literature has created a pressing need for automated survey generation.
+Although recent LLM-based agents have shown promise in automating this task, current progress is hindered by the lack of a standardized, scalable evaluation protocol.
+Existing evaluation methods typically rely on either human evaluation or custom metrics designed to validate specific pipelines, which restricts scalability and hinders fair comparison.
+To address this, we introduce SurGE, a benchmark and evaluation framework tailored for scientific survey generation.
+SurGE provides a large-scale retrieval corpus of over one million papers and expert-validated ground-truth surveys.
+Furthermore, we propose a robust multi-dimensional evaluation protocol that integrates both objective metrics and LLM-based judgments, and empirically verify its high alignment with human experts.
+Our experiments reveal that while agentic pipelines outperform RAG baselines in fluency and structural quality, they still struggle with citation accuracy, highlighting key directions for future research [^1] We have open-sourced all the code and data at: https://github.com/oneal2000/SurGE.
+SurGE: A Benchmark and Evaluation Framework for
+Scientific Survey Generation
 1 Introduction
 The volume of scientific literature has been expanding at an unprecedented rate in recent years.
-For instance, academic archives like arXiv now receive over a thousand new computer science papers daily, more than doubled between 2019 and 2024 (Liang et al., 2025).
-This rapid growth in publications has made the manual creation of comprehensive survey papers increasingly impractical, as the manual collection and synthesis of large volumes of relevant papers is both labor-intensive and time-consuming.
-Faced with this challenge, there is a growing need for automated systems that can generate survey papers effectively.
-While recent advancements in LLM-based agents (Yao et al., 2023; Wang et al., 2024c) offer significant promise for automating this task, their full potential is severely hindered by the lack of reliable, scalable, and standardized evaluation benchmarks.
-Currently, the evaluation of automatically generated surveys largely depends on human-expert reviews, which limit the reproducibility and objectivity of assessments (Tian et al., 2024a).
-Consequently, progress in this field is difficult to quantify, and the relative merits of different methods remain challenging to compare without standardized, multi-faceted evaluation benchmarks.
-To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark that establishes a reproducible standard for the survey generation task, which we formalize as a two-stage process: (1) retrieving relevant papers for a given topic from a large corpus, and (2) synthesizing them into a well-structured survey.
-To support standardized evaluation of both stages, SurGE provides a comprehensive dataset consisting of two key components.
-The first is a large-scale academic corpus containing over one million computer science papers to be used as the search and retrieval pool.
-The second is a collection of test instances, where each instance comprises a research topic (e.g., ”Machine Learning for Information Retrieval”) and its corresponding ground-truth survey.
-Each ground-truth survey is a high-impact, peer-reviewed paper, chosen based on strict citation criteria and subsequently validated for its quality by expert annotators (see Section 3.1), and is provided along with its full list of cited references.
-Beyond dataset construction, another key contribution of this work is a fully automated evaluation framework that enables reproducible and multi-faceted assessment of generated surveys.
-Grounded in established principles for high-quality survey writing (Webster & Watson, 2002; Boote & Beile, 2005; Keele et al., 2007; Pautasso, 2013), our proposed evaluation framework quantifies a survey’s quality across four key dimensions: comprehensiveness, citation accuracy, structural organization, and content quality.
-This automated approach eliminates the reliance on subjective and time-consuming human evaluation, thus establishing a scalable and standardized methodology to guide and measure progress in survey generation.
-To validate the utility of SurGE, we evaluate a range of LLM-based baseline systems on the benchmark.
-These include standard retrieval-augmented generation approaches that first retrieve relevant documents and then generate a survey, as well as more advanced pipelines that incorporate explicit planning (e.g., outline generation) and iterative refinement of the draft.
-The results reveal that even the state-of-the-art systems struggle with the survey generation task.
-For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant citations.
-These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research in this area.
-We believe SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
+Academic archives like arXiv now receive over a thousand new computer science papers daily, a figure that has more than doubled between 2019 and 2024 (Liang et al., 2025).
+This rapid growth in publications has made the manual creation of comprehensive survey papers increasingly impractical, as collecting and synthesizing such vast volumes of information is both labor-intensive and time-consuming.
+Consequently, there is a critical need for automated systems capable of generating high-quality academic surveys.
+Recent LLM-based agents have shown promise in automating this task through hierarchical decomposition or memory-driven mechanisms (Wang et al., 2024; Yan et al., 2025; Lai et al., 2024).
+However, while the feasibility of generation has been established, the field currently lacks a standardized evaluation protocol.
+Existing evaluation methods are often unscalable or strictly tied to specific generation pipelines.
+For example, StepSurvey relies heavily on human evaluation (Tian et al., 2024a), which severely restricts scalability and reproducibility.
+Conversely, framework-oriented works like AutoSurvey and SurveyForge propose custom metrics tailored to validate their own pipelines.
+This "self-validation" paradigm inevitably introduces bias and hinders fair, cross-system comparison.
+As the community moves beyond "making it work" to "making it work well," there is a critical need for a method-agnostic, reproducible, and diagnostic benchmark.
+To address this gap, we introduce SurGE (Survey Generation Evaluation), a comprehensive benchmark and evaluation framework designed to serve as a universal yardstick for the survey generation task.
+We formalize this task as a two-stage process: (1) retrieving relevant papers from a massive corpus, and (2) synthesizing them into a structured survey.
+To support this, SurGE provides a rigorously constructed dataset consisting of a large-scale academic corpus (over one million papers) and a collection of high-quality test instances.
+The ground-truth surveys are sourced from high-impact, peer-reviewed publications, and are also strictly validated by expert annotators to ensure data reliability and ethical compliance (see Section 3.1).
+Beyond dataset construction, a key contribution of SurGE is its robust, multi-dimensional evaluation framework specifically tailored for survey generation.
+Grounded in established principles for high-quality survey writing (Webster and Watson, 2002), we propose a comprehensive evaluation framework that integrates objective, quantifiable indicators (e.g., citation coverage) with LLM-as-a-Judge mechanisms (e.g., assessing logical coherence and content quality).
+Crucially, to ensure the reliability of our proposed evaluation framework, we conduct a rigorous meta-evaluation: we collect fine-grained expert annotations and verify a high correlation between human judgment and our LLM-based metrics.
+This alignment confirms that SurGE serves as a trustworthy and scalable proxy for the evaluation of AI-generated surveys.
+To empirically assess the utility of SurGE, we benchmark diverse LLM-based systems.
+Leveraging the decoupled retrieval-generation design of our framework, we conduct a granular analysis to disentangle whether performance bottlenecks arise from information acquisition or content synthesis.
+Our experiments demonstrate that specialized agentic pipelines consistently outperform standard RAG baselines.
+However, while these agents exhibit superior linguistic fluency, they struggle with information utilization, failing to fully leverage the collected evidence during the generation phase.
 In summary, our contributions are threefold:
   1. 1.
-    We introduce SurGE, a comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
+    We introduce SurGE, a method-agnostic benchmark for survey generation featuring expert-validated ground-truth surveys and a large-scale retrieval corpus.
   2. 2.
@@
   3. 3.
-    We provide extensive baseline results and analyses, offering reference points and highlighting key challenges in this emerging task.
+    We conduct extensive experiments on representative state-of-the-art systems, pointing out specific directions for future optimization.
 2 Task Definition
-We formalize scientific survey generation as a two-stage task.
+We formalize survey generation as a two-stage task.
 Given a topic description $t$ and a large academic corpus $D=\{d_{1},d_{2},\ldots,d_{n}\}$, the goal is to automatically generate a survey article $S$ that provides a structured and comprehensive overview of the topic.
 The process consists of:
-  • Document Collection: A retrieval module or complex agentic retrieval system collects a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
-  • Survey Generation: A generative model composes a well-structured survey $S$ based on the topic $t$ and the retrieved document set $\mathcal{R}_{t}$, including proper citations and a reference list.
-To enable standardized benchmarking of the survey generation task, SurGE offers a comprehensive framework that formalizes the process.
-The benchmark includes 205 expert-selected topics, each paired with a topic description $t$ and a ground-truth survey for evaluation.
-SurGE also provides a large-scale academic corpus $D$ containing over one million papers, which supports the document retrieval stage.
-The generated surveys are then evaluated through an automated system, which quantitatively assesses their quality across four key dimensions.
-Details of the dataset construction process are outlined in Section 3, while the evaluation framework is explained in Section 4.
+  • Information Collection: A retrieval system collects a relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
+  • Survey Generation: A generative system composes a survey $S$ based on the topic $t$ and the retrieved document set $\mathcal{R}_{t}$, including proper citations and a proper reference list.
 3 Dataset Construction
-3.1 Ground-Truth Survey Collection and Expert Annotation
+3.1 Ground-Truth Survey Collection
 To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature.
 Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews.
-To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact (Bornmann & Daniel, 2008); and (iii) the publication date must be between 2020 and 2024.
-Following the initial filtering process, we further refine the SurGE dataset through expert annotation.
-This process aims to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers.
-To this end, we recruited a team of four computer science Ph.D.
-students as annotators.
-Note that annotating a paper did not require a close reading of the entire document.
-Annotators could typically complete the task within 8–9 minutes per paper on average.
+To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact (Bornmann and Daniel, 2008); and (iii) the publication date must be between 2020 and 2024.
+3.2 Expert Annotation
+Building upon the initial automated collection, we implemented a secondary manual filtering phase to ensure the quality of the dataset.
+To this end, we recruited a team of four Ph.D.
+students in computer science, led by a faculty member, to serve as annotators.
 Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) citation quality, which examines the relevance, diversity, and traceability of cited works.
-Each annotator labeled the document as either “usable” or “not usable.” A survey paper was included in the final dataset only if both annotators independently marked it as usable.
-In cases of disagreement, the paper was discarded to maintain a conservative quality threshold.
-To ensure fair and motivated participation, annotators were compensated based on working hours at an average rate of 60 CNY per hour, exceeding the local minimum wage in Beijing.
+Annotation averaged 10 minutes per paper.
+Each annotator labeled the document as either “usable” or “not usable”.
+Papers were retained only if independently labeled as “usable” by both annotators, discarding any disagreements to maintain a conservative quality threshold.
+To ensure motivation, annotators were compensated 60 CNY/hour (much exceeding the local minimum wage).
 Inter-annotator agreement was quantified using Cohen’s Kappa, applied to 250 annotated instances.
-The resulting score of 0.792 indicates substantial agreement and reinforces the reliability of the quality control process.
+The resulting score of 0.792 indicates substantial agreement, reinforcing the reliability of the quality control process.
 After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers.
-The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix F.
-For each of the 205 selected ground-truth surveys, we provide its complete reference list, which serves as the gold standard for evaluating the comprehensiveness of a generated survey.
-The detailed methodology for extracting and processing these references is described in Appendix A.
+The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix H.
 3.3 Academic Corpus Construction
@@
 To ensure adherence to ethical and legal standards, we exclusively collected metadata and did not include full-text PDFs, a practice permitted by arXiv’s Terms of Use, which designates metadata as public domain under the CC0 license (arx, 2025b).
-The corpus was constructed through a two-stage process.
-The initial stage involved seeding the corpus with the references from our 205 ground-truth surveys.
-We systematically retrieved the arXiv metadata for every cited paper that was publicly accessible.
-This process revealed that approximately 30% of the references were unavailable, primarily due to publication in closed-access journals or other restricted venues.
-In the second stage, we expanded the corpus to enhance its comprehensiveness.
+The corpus was constructed through the following process.
+First, we retrieved the arXiv metadata for all publicly accessible papers cited in the ground truth survey.
+This process revealed that around 30% of the references were unavailable, primarily due to publication in closed-access journals or other restricted venues.
+Then we expanded the corpus to enhance its comprehensiveness.
 We queried the official arXiv search API, using keywords and titles from the ground-truth surveys to identify and collect metadata for other topically related papers.
-This methodology resulted in a final retrieval corpus of 1,086,992 unique papers.
+To mitigate topic bias, we augment the corpus with random sample of arXiv computer science papers, drawn independently of the benchmark topics.
+The entire process resulted in a final retrieval corpus of 1,086,992 unique papers.
 For each paper, the corpus provides rich metadata, including its title, authors, abstract, subject categories, publication date, and a direct link to its arXiv page for transparency and verification.
@@
 The resulting SurGE benchmark comprises 205 ground-truth survey papers and a retrieval corpus of 1,086,992 documents.
-Table 1(a) presents the key statistics of our curated dataset.
-To quantitatively analyze the organizational complexity of the surveys, we model the hierarchical section headings (e.g., Section, Subsection) of each paper as a tree structure.
-Our analysis reveals that these surveys are structurally deep, with an average tree depth of 3.07 and a mean of 42.7 nodes (i.e., distinct sections) per document.
-This structural complexity presents a significant challenge for hierarchical text generation.
+Table 1 presents the key statistics of the dataset.
+We model hierarchical headings as trees to quantify the complexity of the ground truth surveys.
+The surveys exhibit deep structures (avg.
+depth 3.07, 42.7 nodes), posing significant challenges for hierarchical text generation.
 Furthermore, the surveys are densely referenced, citing an average of 65.8 papers, which underscores the demand for high-recall information collection.
-Table 1(b) details the pre-processed fields for each survey instance, which include not only standard metadata but also the parsed structural tree and the ground-truth list of cited documents, enabling a fine-grained, multi-faceted evaluation of system-generated surveys.
+Detailed descriptions of the metadata fields for both the survey instances and the retrieval corpus are provided in Appendix B.
 3.5 Ethical Considerations and Licensing
@@
 This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata (arx, 2025a).
-We have prioritized transparency, reproducibility, and ethical considerations throughout dataset construction.
-To support open science, we have publicly released the SurGE dataset, accompanying metadata, and all associated processing scripts on our official GitHub repository [^2] https://github.com/oneal2000/SurGE.
+A more detailed discussion of ethical considerations is provided in Section §10.
+To support open science, we have publicly released the SurGE dataset, accompanying metadata, and all associated processing scripts on our official GitHub repository.
 The dataset and codebase are distributed under the MIT license, granting researchers and developers unrestricted access and modification rights.
-Regular updates will ensure continued relevance and alignment with evolving research trends and standards.
-Figure 1: Overview of the SurGE Benchmark. (a) Summary statistics of the curated survey dataset and its associated retrieval corpus. (b) Metadata of the pre-processed survey dataset used in SurGE.
-(a) Basic Statistics of the SurGE Dataset and Corpus
+Table 1: Key statistics of the SurGE benchmark dataset and its associated retrieval corpus.
 Statistic Number
@@
 Average Citations per Paper 65.78
-Average Citations per Section 1.577
 Corpus Size 1,086,992
 Average Abstract Length (words) 156.57
-(b) Fields in the Pre-processed Survey Dataset
-Field Description
-SurveyID A unique identifier for the survey.
-Authors List of contributing researchers.
-Title The title of the survey paper.
-Year The publication year of the survey.
-Date The timestamp of publication.
-Category Subject classification following the arXiv taxonomy.
-Abstract The abstract of the survey paper.
-Structure Hierarchical representation of the survey.
-All_Cites List of document IDs cited in the survey.
 4 Evaluation Framework
@@
 $\mathcal{R}=\frac{|{R}_{GT}\cap{R}_{G}|}{|{R}_{GT}|},$ (1)
-While the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited, peer-reviewed publications (detailed in §3.1).
+Crucially, while the GT reference set is not assumed to be perfectly complete, it serves as the best available proxy for expert consensus on a topic’s core literature, given that our GT surveys are highly cited and peer-reviewed.
 We therefore interpret this metric not as a measure of absolute completeness, but as a pragmatic metric for evaluating a system’s ability to identify the central body of work validated by the research community.
@@
 Citation accuracy is another critical aspect of a high-quality survey.
-Each citation in a survey should be thematically relevant to the overall topic of the survey, and it must be contextually appropriate in terms of both the section and sentence in which it appears.
-To evaluate this aspect, we introduce the metric of Citation Accuracy, which evaluates each citation in the survey across three levels.
+Each citation in a survey should be thematically relevant to the overall topic, and it must be contextually appropriate in terms of both the section and sentence in which it appears.
+To evaluate this aspect, we introduce Citation Accuracy, which evaluates each citation in the survey across three levels.
 First, at the document level, we assess whether a cited paper is thematically relevant to the overall topic of the survey.
@@
 Finally, at the sentence level, we verify whether a citation supports the specific claim made in the sentence where it is cited.
-To automate this evaluation, we employ a Natural Language Inference (NLI) model (nli-deberta-v3-base [^3] https://huggingface.co/cross-encoder/nli-deberta-v3-base) to assess the relevance of each citation.
+To automate this evaluation, we employ a Natural Language Inference (NLI) model (nli-deberta-v3-base He et al.
+(2021) to assess the relevance of each citation.
 An NLI model is designed to determine the logical relationship between two text snippets: a premise and a hypothesis.
-The model then predicts the relationship between these two components, providing probabilities for the following labels: Entailment, Neutral, Contradiction.
-Due to its ability to capture semantic relationships, NLI has become a standard method for evaluating the factual consistency and relevance of text generated by LLMs, rendering it an ideal tool for our task.
+The model predicts the relationship between these two components, providing probabilities for the following labels: Entailment, Neutral, Contradiction.
+Due to its ability to capture semantic relationships, NLI has become a standard method for evaluating the factual consistency of text generated by LLMs Min et al.
+(2023).
 For our specific evaluation, we implement this three-level check by framing it as a series of NLI tasks.
@@
 Conversely, any citation $r$ that is present in the ground-truth survey’s bibliography is assigned a document-level score of $y_{d}(r)=1$.
-For all other cases, the relevance score $y_{x}(r)$ is determined by the NLI model’s output probabilities for the labels Entailment, Neutral, Contradiction.
-The probabilities are mapped to a final score for each unit as follows:
+For the remaining cases, the relevance score $y_{x}(r)$ is derived from the NLI model’s output distribution.
+Let $p_{\text{ent}}$, $p_{\text{neu}}$, and $p_{\text{con}}$ denote the predicted probabilities for the Entailment, Neutral, and Contradiction labels, respectively.
+We map these probabilities to a discrete score as follows:
 […]
+$(\text{where }x\in\{d,s,t\})$
 Finally, we aggregate the individual citation scores to produce three final metrics for the survey: Document-level Accuracy ($R_{d}$), Section-level Accuracy ($R_{s}$), and Sentence-level Accuracy ($R_{t}$).
@@
 Our first metric, the Structure Quality Score (SQS), addresses the high-level organization.
-It holistically assesses the alignment between the generated and ground-truth outlines by comparing their overall structure, semantic coherence, and topical progression.
+It assesses the alignment between the generated and ground-truth outlines by comparing the similarity of their headings in terms of structure, meaning, and wording.
 Complementing this, our second metric, Soft-Heading Recall (SHR), provides a fine-grained evaluation of heading alignment.
@@
 Structure Quality Score (SQS).
-SQS evaluates the overall quality of a generated survey’s structure based on the hierarchical list of its section headings.
-To compute this score, we adopt the LLM-as-a-Judge paradigm, leveraging GPT-4o as the evaluator.
-Specifically, we provide the LLM with both the generated and ground-truth outlines and prompt it to assign a quality score.
-To guide the LLM’s evaluation, we have carefully designed a detailed instruction prompt that includes a comprehensive scoring rubric on a scale from 0 to 5.
-The complete prompt is shown in Appendix E.
+SQS evaluates the structural fidelity of a generated survey by assessing its alignment with the ground truth.
+To capture both the semantic content and the logical organization of the sections, we adopt the LLM-as-a-Judge paradigm.
+Specifically, we provide the LLM with the hierarchical heading structures from both the generated and ground-truth surveys.
+Crucially, to preserve the structural hierarchy (e.g., section vs.
+subsection), we formulate these inputs using explicit markup (e.g., Markdown or LaTeX) rather than flat lists.
+The model is then prompted to assign a similarity score based on the correspondence of structure, meaning, and wording, following a comprehensive scoring rubric (see Appendix G).
 Soft-Heading Recall (SHR).
@@
 Therefore, we define the soft intersection cardinality using the inclusion-exclusion principle:
-$\mathcal{S}(H_{P}\cap H_{GT})=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})-\mathcal{S}(H_{P}\cup H_{GT}).$ (6)
+$\begin{split}\mathcal{S}(H_{P}\cap H_{GT})&=\mathcal{S}(H_{P})+\mathcal{S}(H_{GT})\\ &\quad-\mathcal{S}(H_{P}\cup H_{GT}).\end{split}$} (6)
 The core idea lies in the union term, $\mathcal{S}(H_{P}\cup H_{GT}$.
@@
 The evaluation is based on five criteria: fluency and coherence, logical clarity, avoidance of redundancy, clarity of description, and absence of errors.
-To guide the LLM’s evaluation, we designed a detailed instruction prompt for the LLM, which is provided in the Appendix E.
+To guide the LLM’s evaluation, we designed a detailed instruction prompt for the LLM, which is provided in the Appendix G.
 Each section is scored on a scale of 0 to 5, where a higher score reflects superior fluency, logical progression, and clarity.
 Considering the context length limitations of the LLM, we have it score each survey section by section, and the final score is the average of the scores from all sections.
-As supplementary measures, we also compute ROUGE and BLEU scores, which quantify n-gram overlap between the generated and ground-truth surveys.
-While these are well-established metrics in text generation, their role is secondary in our framework, serving as additional checks for content fidelity rather than a primary assessment method.
+Additionally, we also report ROUGE and BLEU scores in our main experiments for completeness.
+However, both ROUGE and BLEU are provided strictly for reference and do not constitute part of the metric suite in our SurGE benchmark.
+Traditional lexical matching metrics fall short in accurately measuring the performance of open-ended scientific generation, as they fail to capture the semantic correctness and logical flow essential for high-quality surveys.
+4.5 Meta-Evaluation: Human Alignment
+Alignment with expert consensus is a prerequisite for automated evaluation. To validate our metrics, we compared the rankings produced by our evaluator against human expert judgments.
+Setup.
+We selected a representative subset of baselines (Section 5.2): RAG, StepSurvey, and AutoSurvey, all utilizing the Qwen2.5-14B-Instruct backbone.
+The human evaluation was conducted by doctoral researchers in Computer Science, all of whom possess prior experience in academic peer review.
+For each test query, these experts ranked four anonymous surveys (the three LLM-outputs plus the human-written Ground Truth) and were tasked with producing two independent ranking lists based on: (1) Structural Quality, adhering to the definitions in Section 4.3; and (2) Content Quality, following the metrics in Section 4.4.
+We quantified alignment using Kendall’s $\tau$ (Kendall, 1938) and Spearman’s $\rho$ (Spearman, 1904).
+As human-written surveys significantly outperform models while model-to-model differences are more subtle, we evaluate alignment in two distinct settings: (1) w/ GT: calculating correlation across the full list to verify the validity of our framework in identifying the gold standard; and (2) w/o GT: calculating correlation excluding the GT to assess the discriminative power among competitive baselines.
+Results.
+Table 2 presents the human-model alignment measured by Kendall’s $\tau$ and Spearman’s $\rho$.
+In the w/ GT setting, SurGE achieves a high $\tau$ of $0.805$, confirming strong alignment in distinguishing human-level scholarship from machine-generated text.
+Crucially, in the more challenging w/o GT setting, our framework maintains a substantial correlation (Spearman’s $0.707$, Kendall’s $0.610$).
+Prior studies in NLG evaluation suggest that these values indicate strong reliability.
+For instance, meta-evaluations of summarization metrics often report Spearman correlations in the range of $0.5$ to $0.7$ as state-of-the-art performance (Liu et al., 2023; Fabbri et al., 2021).
+Table 2: Human-LLM alignment meta-evaluation.
+We measure the correlation between human rankings and our metrics: Structure (SQS) and Content (CQS).
+w/ GT includes human-written ground truth in the ranking list, while w/o GT considers generated baselines only.
+Metric SQS CQS
+w/ GT w/o GT w/ GT w/o GT
+Kendall’s $\tau$ ($\uparrow$) 0.805 0.610 0.797 0.594
+Spearman’s $\rho$ ($\uparrow$) 0.883 0.707 0.878 0.695
 5 Experimental Setup
@@
 Each baseline follows a two-stage pipeline: (1) retrieving a set of potentially relevant papers for a given topic, and (2) organizing and summarizing the retrieved papers to produce a structured survey.
+5.1 Paper Retriever
 For fair comparison, all baselines share the same dense retriever for the first stage.
-In the following subsections, we first describe the training of the shared Paper Retriever, followed by the baseline selection and implementation details.
-5.1 Paper Retriever Training
-We employ a dual-encoder architecture for retrieval, initialized with roberta-base.
-The training process leverages the benchmark dataset, where each topic description $t$ (introduced in §2) serves as the query $q$.
-For a query $q$ and a paper abstract $d$, we construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$.
-Formally, let $X(q)=[\mathrm{CLS}]\,q\,[\mathrm{SEP}]$ and $X(d)=[\mathrm{CLS}]\,d\,[\mathrm{SEP}]$.
-We then feed these tokens into roberta-base to obtain the contextualized embedding of the $\mathrm{CLS}]$ token:
-$\mathrm{Emb}(X)=\mathrm{transformer}_{[\mathrm{CLS}]}(X).$ (7)
-The similarity score between the $q$ and the $d$ is computed as the dot product of their embeddings:
-$S(q,d)=\mathrm{Emb}(X(q))^{\top}\cdot\mathrm{Emb}(X(d)).$ (8)
-During training, each query $Q$ is paired with the relevant documents $d^{+}$ from the ground truth paper to form the positive samples, while negative samples $d^{-}\in N$ are randomly sampled from the corpus.
-The retriever is optimized via the softmax cross-entropy loss:
-$\mathcal{L}(Q,d^{+},N)=-\log\frac{\exp\bigl(S(Q,d^{+})\bigr)}{\exp\bigl(S(Q,d^{+})\bigr)+\sum_{d^{-}\in N}\exp\bigl(S(Q,d^{-})\bigr)}.$ (9)
-This objective encourages the model to assign higher scores to relevant papers while minimizing scores for irrelevant ones.
-After training, we use the trained retriever to retrieve the top-ranked papers for each query, thereby providing a collection of relevant papers for the subsequent generation stage.
+We employ a dual-encoder architecture initialized with roberta-base Liu et al.
+(2019).
+The retriever is trained using a contrastive learning objective, where the topic description serves as the query and the paper abstract as the document.
+Specifically, we optimize the model via softmax cross-entropy loss to maximize the dot-product similarity between queries and their ground-truth relevant papers while distinguishing them from negative samples.
+Further details on the mathematical formulation and training setup can be found in Appendix D.
 5.2 Baselines
-We selected the following three survey generation baseline methods for our experiments. Detailed descriptions of each method are provided in the appendices.
-  • Retrieval-Augmented Generation (RAG): This straightforward approach retrieves relevant papers, summarizes them in chunks, and then merges these summaries to form the final survey.
-A detailed description is provided in Appendix D.1.
-  • AutoSurvey (Wang et al., 2024d): This method employs a multi-stage, outline-driven pipeline.
+We select four representative survey generation methods as baselines, with detailed descriptions and implementations provided in the appendices.
+  • Retrieval-Augmented Generation Lewis et al.
+(2020).
+RAG approach retrieves relevant papers, summarizes them in chunks, and merges these summaries to form the final survey.
+A detailed description is provided in Appendix E.1.
+  • AutoSurvey (Wang et al., 2024) employs a multi-stage, outline-driven pipeline.
 It first generates a high-level outline from retrieved papers and then iteratively expands and refines each section.
-Further details can be found in Appendix D.2.
-  • StepSurvey (Lai et al., 2024b): This approach adopts a granular, step-by-step process.
-It starts by generating a title and primary headings, which then guide the incremental drafting of each subsection.
-The methodology is described in detail in Appendix D.3.
-5.3 Implementation Details.
-For the training of Paper Retriever, we randomly split the dataset into a training set and a test set at a ratio of 4:1.
-We adopt the AdamW optimizer for model optimization, the learning rate is set to $5\times 10^{-6}$, and the epoch is set to $10$.
-During the training process, we adopt mixed-precision (fp16) training.
-At inference time, each query retrieves the top 100 relevant papers according to the similarity score.
+Further details can be found in Appendix E.2.
+  • StepSurvey (Lai et al., 2024) adopts a step-by-step generation process, starting by generating a title and primary headings, which then guide the incremental drafting of each subsection.
+The methodology is detailed in Appendix E.3.
+  • SurveyForge (Yan et al., 2025) introduces a two-stage pipeline comprising heuristic outline generation and memory-driven content writing. Detailed information is provided in Appendix E.4.
+5.3 Implementation Details
+Retriever Training.
+For the training of the Paper Retriever, we randomly split the dataset into training and test sets at a ratio of 4:1.
 The retriever is initialized using the pre-trained RoBERTa model (Liu et al., 2019).
-To ensure a fair comparison, we utilize the Qwen2.5-14B-Instruct model (Yang et al., 2024a) as the base LLM for all baseline methods.
-For the generation configuration of LLMs, all experiments are conducted using the publicly available implementations provided by Hugging Face.
-We utilize the default hyperparameters and the chat template as outlined in the official Hugging Face repository  [^4] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
-All experiments are conducted on a GPU server with 1TB RAM and eight NVIDIA A100 GPUs, each with 40GB of memory.
+We adopt the AdamW optimizer with a learning rate of $5\times 10^{-6}$, and the model is trained for $10$ epochs using mixed-precision (fp16) training.
+At inference time, each query retrieves the top-100 relevant papers based on similarity scores.
+Model Configuration.
+To ensure a fair comparison among the pipeline-based baseline methods (AutoSurvey, StepSurvey, and SurveyForge), which originally utilize different underlying models, we unify their base LLM to Qwen2.5-14B-Instruct (Yang et al., 2024).
+For the Retrieval-Augmented Generation (RAG) baseline, we extend our evaluation to include three distinct backbone models to analyze performance across different model capabilities: Qwen2.5-14B-Instruct, GPT-4o Hurst et al.
+(2024), and Gemini-2.5-Pro-Thinking Comanici et al.
+(2025).
+For the generation configuration of open-source models, we utilize the implementation provided by Hugging Face with default hyperparameters and the official chat template [^2] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct.
+For the embedding model of the SHR metric, we use the bge-large-en-v1.5 model Chen et al.
+(2024).
+All experiments are conducted on a GPU server equipped with 1TB RAM and eight NVIDIA A100 GPUs (40GB memory each).
 6 Experimental Results
-This section presents the detailed evaluation of the selected survey generation baselines.
-We first evaluate the performance of the shared Paper Retriever (§6.1).
-This analysis is crucial as it establishes the theoretical upper-bound performance for the comprehensiveness metric and highlights the challenges posed by the initial retrieval stage.
-Following this, we assess the complete end-to-end performance of the three baselines (§6.2) based on our proposed evaluation framework.
-Table 1: Comparison of retrieval models on recalling ground-truth cited papers. The metric is Recall@k, where $k$ is the number of top documents retrieved. Best results are in bold.
-Model Recall@20 Recall@30 Recall@100 Recall@200 Recall@500 Recall@1000
+This section presents the experimental results. We begin by analyzing the performance of the Paper Retriever (§6.1), then analysis the end-to-end survey generation performance of the baselines (§6.2).
+6.1 Evaluation of the Retrieval Stage
+A crucial question in our two-stage pipeline is whether the low reference recall of baseline systems stems from the retriever’s inability to find relevant papers (a data coverage issue) or the generator’s inability to identify them (a capability issue).
+To disentangle these factors and validate the quality of our constructed corpus, we evaluate the performance of our fine-tuned dense retriever in isolation.
+We compare our dense retriever (PR) against the lexical baseline BM25 (Robertson et al., 2009), using Recall@k.
+As shown in Table 3, our fine-tuned Paper Retriever substantially outperforms the BM25 baseline.
+More importantly, the results provide strong evidence for the validity of our benchmark data.
+At $k=1000$, the retriever successfully recalls 68.05% of the ground-truth references.
+This high upper-bound coverage confirms that the vast majority of expert-selected references are indeed present and retrievable within our corpus, addressing potential concerns regarding the completeness of the data collection process.
+Furthermore, the substantial disparity between the retrieval upper bound (68.05%) and end-to-end baseline performance ($10\%$) indicates that the bottleneck of survey generation lies in the agents’ capacity to select and synthesize information.
+While this confirms the value of SurGE as a challenging testbed, the retrieval stage itself presents further opportunities for optimization.
+Future research could investigate sophisticated retrieval paradigms such as task-oriented search agents.
+Table 3: Comparison of retrieval models on recalling ground-truth cited papers. The metric is Recall@k (R@k), where $k$ is the number of top documents retrieved. Best results are in bold.
+Model R@20 R@30 R@100 R@200 R@500 R@1000
 BM25 0.0548 0.0652 0.1193 0.1596 0.2213 0.2715
-Paper Retriever 0.1706 0.2145 0.3665 0.4681 0.6011 0.6805
-6.1 Analysis of Retrieval Performance
-A crucial question in our two-stage pipeline is whether performance limitations stem from the retriever’s inability to find relevant papers or the generator’s inability to use them.
-To disentangle these factors and quantify the retrieval bottleneck, we evaluate the performance of our fine-tuned dense retriever in isolation.
-This analysis establishes the theoretical upper bound for the reference coverage that our end-to-end systems can achieve.
-We compare our dense retriever against the lexical baseline BM25 (Robertson et al., 2009), using Recall@k as the evaluation metric.
-This metric measures the percentage of ground-truth papers from the reference survey that are present in the top-k retrieved documents.
-As shown in Table 1, our fine-tuned Paper Retriever substantially outperforms the BM25 baseline across all values of $k$.
-The performance gap underscores the inadequacy of lexical search, which struggles to find semantically relevant papers that may not share overlapping keywords.
-On the other hand, our dense retriever can capture deeper semantic relationships, thus it is far more effective than BM25.
-However, the results also reveal a critical bottleneck in the survey generation pipeline.
-Even when retrieving the top 1000 documents (k=1000), Paper Retriever’s recall reaches only 68.05% of the ground-truth papers.
-This performance ceiling imposes a hard upper bound on the downstream generator, as it cannot synthesize information from papers it never receives.
-Consequently, this performance ceiling underscores that the survey generation task needs more sophisticated retrieval paradigms, such as employing search agents or multi-agent systems powered by large language models (Zhang et al., 2024; Tang et al., 2025; Li et al., 2025).
-Table 2: Main experimental results comparing different survey generation baselines across four dimensions: Comprehensiveness (Comp.), Citation Accuracy, Structural Quality, and Content Quality.
+PR 0.1706 0.2145 0.3665 0.4681 0.6011 0.6805
+Table 4: Experimental results comparing survey generation baselines across four dimensions: Comprehensiveness (Comp.), Citation Accuracy, Structural Quality, and Content Quality.
 Metrics include Recall, Document/Section/Sentence-level Citation Accuracy (Doc-Acc, Sec-Acc, Sent-Acc), Structure Quality Score (SQS), Soft-Heading Recall (SHR), ROUGE-L, BLEU, and Content Quality Score (CQS).
-The best results are in bold, and the second-best results are underlined.
+The best results within each group are in bold.
 Comp. Citation Accuracy Structural Quality Content Quality
 Baseline Recall Doc-Acc Sec-Acc Sent-Acc SQS SHR R-L BLEU CQS
-RAG 0.0214 0.2857 0.2502 0.2500 0.6829 0.7900 0.1519 10.38 4.6723
+General-Purpose RAG Baselines
+RAG-Qwen 0.0214 0.2857 0.2502 0.2500 0.6829 0.7900 0.1519 10.38 4.6723
+RAG-GPT 0.0419 0.4525 0.3166 0.3226 0.7561 0.6659 0.1409 8.36 4.9250
+RAG-Gemini 0.0784 0.4680 0.4056 0.4124 1.2927 0.7591 0.1442 8.37 4.5619
+Specialized Survey Generation Agents
 AutoSurvey 0.0351 0.3617 0.4935 0.4870 1.3902 0.9697 0.1578 10.44 4.7390
-StepSruvey 0.0630 0.4576 0.4571 0.4636 1.1951 0.9763 0.1590 12.02 4.8451
-6.2 Overall Performance of Survey Generation Baselines
-Our evaluation now transitions from retrieval performance to the end-to-end survey generation task.
-While the retriever supplies the models with a substantial portion (36.65%) of the ground-truth references, a detailed analysis of the final surveys reveals a significant bottleneck in the generation stage.
-Table 2 presents a multi-faceted comparison of our three baselines across the four dimensions of survey quality defined by our benchmark.
-This comprehensive evaluation yields several key insights:
-(1) The Generation Stage is the Primary Bottleneck.
-A striking finding is the dramatic performance drop-off between paper retrieval and final survey synthesis.
-Despite the retriever providing access to 36.65% of ground-truth references (Recall@100), the best-performing model, StepSurvey, incorporates only 6.30% of them into its final output.
-The standard RAG baseline fares even worse, with a final recall of just 2.14%.
-This stark disparity highlights that the core challenge lies not in finding relevant papers, but in the generator’s ability to effectively identify and integrate crucial information from the retrieved set.
-Nevertheless, the results also affirm the value of structured pipelines, as AutoSurvey (3.51%) and StepSurvey (6.30%) significantly outperform the naive RAG approach in comprehensiveness.
-(2) AutoSurvey Excels in Local Coherence and Structural Quality.
-AutoSurvey demonstrates particular strength in producing locally coherent and well-organized content.
-As shown in Table 2, it achieves the highest scores for both Section-level Citation Accuracy (Sec-Acc: 0.4935) and Sentence-level Citation Accuracy (Sent-Acc: 0.4870).
-This suggests its iterative, section-by-section refinement process is highly effective at placing citations within their correct immediate context.
-Furthermore, their planning approach results in the highest Structure Quality Score (SQS) of 1.3902, indicating superior hierarchical organization in the generated survey.
-(3) StepSurvey Achieves Superior Coverage and Global Content Quality.
-In contrast, StepSurvey’s strengths lie in metrics related to holistic coverage and overall content quality.
-It attains the highest Recall (0.0630) and Document-level Citation Accuracy (Doc-Acc: 0.4576), demonstrating a better ability to cover the breadth of the topic and align cited works with the survey’s main theme.
-Its multi-phase workflow that generates headings, then subtopics, and finally assigns citations leads to the best content quality.
-This is evidenced by its top scores in text similarity metrics (ROUGE-L: 0.1590, BLEU: 12.02) as well as the highest Content Quality Score (CQS) of 4.8451, which reflects stronger logical flow and presentation.
-(4) A Trade-off Between Local Precision and Global Coverage.
-The distinct strengths of AutoSurvey and StepSurvey reveal a fundamental trade-off in planning strategies.
-AutoSurvey’s iterative refinement of individual sections fosters high-quality local structure and precise, fine-grained citation placement.
-Conversely, StepSurvey’s hierarchical, topic-first approach yields better overall topic coverage and more coherent, globally relevant content.
-While both advanced methods significantly outperform the standard RAG pipeline, neither excels across all dimensions.
-In conclusion, our proposed SurGE benchmark effectively diagnoses the weaknesses of current survey generation systems.
-The results clearly indicate that advanced, multi-stage planning is crucial, but a substantial gap remains between machine-generated and expert-written surveys, particularly in comprehensiveness.
-The observed trade-off between local and global optimization strategies highlights a key challenge for future research: developing hybrid models that can combine the fine-grained accuracy of AutoSurvey with the broad thematic coverage and coherence of StepSurvey.
+StepSurvey 0.0630 0.4576 0.4571 0.4636 1.1951 0.9763 0.1590 12.02 4.8451
+SurveyForge 0.0868 0.4719 0.4651 0.4772 1.0537 0.9493 0.1443 8.99 4.8070
+6.2 Evaluation of End-to-End Generation
+Table 4 presents the comparative results across four critical dimensions.
+Our analysis reveals four key observations regarding the trade-offs between retrieval capabilities, structural coherence, and content fidelity: (1) Surface-level fluency does not imply factual accuracy.
+We observe that RAG-Qwen achieves competitive scores in n-gram metrics like ROUGE-L and BLEU, outperforming the more sophisticated RAG-GPT.
+However, it suffers from the lowest Citation Accuracy and Recall.
+This disparity highlights a "fluency hallucination" trap, where naive RAG models can generate coherent but factually ungrounded text.
+This underscores the necessity of relying on fact-centric metrics (e.g., Doc-Acc, CQS) rather than traditional n-gram metrics for evaluating scientific survey generation.
+(2) Agentic workflows are essential for structural quality.
+Methods incorporating explicit planning stages (AutoSurvey, StepSurvey, and SurveyForge) consistently outperform standard RAG approaches in maintaining structural integrity.
+This confirms that decomposing the generation process into outline planning and content filling is critical for capturing the hierarchical logical structures characteristic of academic surveys.
+(3) Specialized mechanisms compensate for limited model capacity.
+Comparing RAG variants demonstrates that model capability matters, as RAG-Gemini significantly outperforms RAG-Qwen.
+However, SurveyForge reverses this reliance on scale: by utilizing Qwen-14B, it not only outperforms its direct counterpart RAG-Qwen, but also surpasses RAG-Gemini in Citation Accuracy.
+This suggests that a well-structured agentic framework can effectively unlock the potential of smaller models, enabling them to outperform larger baselines in complex, reference-dependent generation tasks.
+(4) Different agentic frameworks yield divergent strengths.
+While SurveyForge excels in retrieving accurate literature (Recall and Doc-Acc), StepSurvey dominates in content quality metrics, achieving the highest CQS.
+This indicates that SurveyForge’s memory-driven mechanism ensures better coverage of relevant papers, whereas StepSurvey’s granular, step-by-step drafting process yields superior linguistic quality and coherence.
 7 Related Work
-7.1 Retrieval-Augmented Generation
-Large Language Models (LLMs) are inherently limited by their static, pre-trained parametric knowledge.
-To address these limitations, Retrieval-Augmented Generation (RAG) has emerged as a key paradigm (Gao et al., 2023; Lewis et al., 2020; Dong et al., 2025; Tu et al., 2025; Su et al., 2025a).
-By grounding the model in external knowledge, RAG directly addresses several fundamental limitations of LLMs, offering a robust mechanism to mitigate hallucinations (Ji et al., 2023; Su et al., 2024d; b; Wang et al., 2025a), facilitate knowledge updating (Fang et al., 2024a; Wang et al., 2024b; a; 2025b), and enable effective domain adaptation (Yang et al., 2024b; Su et al., 2025c; 2024a; 2024e).
-The conventional approach to traditional RAG is built upon the ”Retrieval-then-Read” paradigm (Borgeaud et al., 2022; Guu et al., 2020; Lewis et al., 2020).
-Within this framework, a user’s query triggers a search module for relevant documents within a large-scale external corpus.
-This retrieval step is carried out by either an external retriever (Zhai, 2008; Su et al., 2023b; Robertson et al., 2009; Su et al., 2023a; Ma et al., 2023; Fang et al., 2024b) or a more sophisticated retrieval system (Su et al., 2023c; Salemi & Zamani, 2024; Chen et al., 2022).
-Building upon this foundation, recent work has proposed more advanced RAG architectures to improve efficiency and effectiveness.
-For instance, Dynamic RAG (Jiang et al., 2022; Su et al., 2024c; Yao et al., 2024) moves beyond a single retrieval step by adaptively triggering the retriever during generation, specifically when the LLM is uncertain during the generation process.
-From another angle, GraphRAG (Edge et al., 2024) enhances the knowledge source by querying pre-constructed knowledge graphs instead of unstructured text, allowing it to retrieve interconnected facts and relationships.
-Furthermore, the Parametric RAG paradigm  (Su et al., 2025b; Tan et al., 2025; Fleshman & Van Durme, 2025) alters the knowledge injection step by directly injecting retrieved knowledge into the LLM’s parameters.
-The scientific survey generation task, which is the focus of our SurGE benchmark, presents a significant challenge for even these advanced RAG systems.
-Unlike typical question-answering, survey generation demands the synthesis of a large, diverse set of documents into a coherent, well-structured survey paper.
-Therefore, while RAG provides the foundational technology, our SurGE benchmark is specifically designed to push the boundaries of current models by rigorously evaluating their capabilities in large-scale multi-document synthesis and structured content creation.
-7.2 Long-Form Text Generation and Evaluation
-Long-form text generation is substantially more challenging than short-text generation due to its inherent requirements for sustained coherence and rich contextual understanding.
-Early approaches mainly used generative adversarial networks and reinforcement learning to conduct long-sequence generation (Guo et al., 2018).
-More recently, large language models have emerged as a strong tool for this task, offering advanced capabilities to handle long-text generation.
-For example, structured planning techniques and specialized inference mechanisms are proposed (Sloan et al., 2024; Jin et al., 2024) to generate consistent and high-quality clinical reports.
-Similarly, hierarchical planning frameworks have demonstrated that content control and multi-constraint instruction following can significantly enhance logical flow and topic coverage (Hu et al., 2022; Pham et al., 2024).
-Beyond medical or other task-specific applications, context-driven retrieval strategies, such as tree-structured retrieval, can support open-domain long-text generation by guiding the model through extensive knowledge sources (Roy et al., 2024).
-The effective evaluation framework is vital for measuring the quality, factualness, and user-centric utility of the long-form text generation task.
-Traditional metrics, designed for shorter texts, often fail to capture the intricacies of longer outputs.
-Recent work has introduced task-focused benchmarks that emphasize user-oriented objectives, such as personalized writing or domain-specific content generation (Kumar et al., 2024; Salemi et al., 2025).
-In parallel, factuality assessment has attracted growing interest, with methods proposed to evaluate both verifiable and unverifiable claims.
-Metrics such as VERISCORE and FACTSCORE break down generated text into atomic facts, checking each for consistency against reliable sources (Song et al., 2024; Min et al., 2023).
-Beyond factual correctness, coherence and structural quality have been studied extensively.
-Benchmarks like LongGenBench and HelloBench underscore the importance of evaluating a model’s ability to maintain logical organization and clarity over extended passages (Wu et al., 2025; Que et al., 2024).
-7.3 Survey Generation
-In the domain of scientific writing, survey generation involves distilling extensive textual resources into a coherent and structured overview.
-Recent advances in AI-assisted systems have provided prompting-based approaches to expedite the drafting process while preserving content accuracy (Gero et al., 2022; Kacena et al., 2024).
-One of the most commonly used approaches is retrieval-augmented generation, which combines large-scale knowledge retrieved from documents with language generation empowered by LLMs to yield factually comprehensive overviews (Lewis et al., 2021).
-Retrieval-augmented generation is often initiated with dense retrieval methods based on dual-encoder architectures to identify highly relevant documents (Karpukhin et al., 2020).
-Once these documents are retrieved, summarization techniques—spanning top-down, bottom-up, and graph-based ranking methods—play a pivotal role in producing concise yet faithful summaries (Nayeem & Rafiei, 2024; Pang et al., 2022; Bleiweiss, 2023).
-Building on these retrieval and summarization-based methodologies, automated literature survey generation has garnered increasing attention (Tian et al., 2024a; Lai et al., 2024a).
-However, existing techniques depend on limited ground truths and employ coarse evaluation metrics, resulting in oversimplified assessments of survey quality (Wang et al., 2024d).
-To address these challenges, we present a refined ground truth and a multi-dimensional evaluation framework that emphasizes both accuracy and structural coherence.
-By evaluating quality through multiple dimensions, our proposed framework advances the capabilities of automated survey generation, offering a more comprehensive and rigorous approach to summarizing scientific literature.
+Recent advancements in LLMs have shifted the paradigm from simple summarization to full-scale survey generation.
+Several pioneering works have demonstrated the feasibility of this task through agentic pipelines.
+For instance, AutoSurvey (Wang et al., 2024) and StepSurvey (Lai et al., 2024) tackle the long-context challenge via hierarchical decomposition, employing outline-driven or step-by-step strategies to refine content from headings to subsections iteratively.
+To enhance information utilization, SurveyForge (Yan et al., 2025) integrates a Scholar Navigation Agent (SANA) with memory-driven writing mechanisms, while SurveyX (Liang et al., 2025) introduces an AttributeTree structure to organize reference data before the generation phase.
+Collectively, these works confirm that LLM-based agentic systems possess the capability to create complex academic surveys.
+However, a standardized evaluation protocol remains absent, as existing assessments rely heavily on unscalable human judgments or biased self-validation protocols.
+SurGE addresses this critical gap by establishing the first method-agnostic benchmark to rigorously measure progress and pinpoint bottlenecks in both retrieval and generation stages.
+Due to space constraints, we provide a more comprehensive discussion of related work in Appendix F.
 8 Conclusion
-In this paper, we introduce SurGE, a comprehensive benchmark designed to address the critical need for standardized and reproducible evaluation in automated scientific survey generation.
+In this paper, we introduce SurGE, a benchmark and evaluation framework designed to address the critical need for standardized and reproducible evaluation in automated scientific survey generation.
 SurGE provides a large-scale academic corpus, a set of expert-written ground-truth surveys, and a fully automated framework to evaluate surveys on comprehensiveness, citation accuracy, structure, and content.
-Our experiments reveal significant limitations in state-of-the-art LLM-based systems, highlighting challenges such as incomplete topic coverage and reference hallucination.
-We believe SurGE will catalyze future research and guide the development of more effective LLM-based systems at the intersection of information retrieval and generative AI for this important task.
-9 Ethics Statement
-The SurGE benchmark aims to facilitate the advancement of automated survey generation and evaluation in the context of scientific research.
-By utilizing publicly available metadata from arXiv, our work prioritizes transparency and open access.
-We do not host or redistribute arXiv’s PDF content, ensuring compliance with arXiv’s Terms of Use and the CC0 public-domain dedication for metadata.
-While the use of large-scale academic data in SurGE poses minimal ethical risks, we emphasize the importance of ethical data usage, privacy, and responsible development of AI systems.
-We encourage the research community to consider these aspects when applying or building upon SurGE.
-10 Reproducibility Statement
-To ensure reproducibility, all code, data, and models for the SurGE benchmark are open-sourced and available at our official GitHub Repository [^5] https://github.com/oneal2000/SurGE.
-Detailed descriptions of the dataset construction, processing steps, and evaluation methods can be found at our official GitHub Repository and appendices.
-Our open-access approach is designed to facilitate the replication of results and encourage future research in the field of survey generation using large language models.
+Our experiments reveal the limitations in SOTA LLM-based survey generation systems, highlighting challenges such as incomplete topic coverage and reference hallucination.
+We believe SurGE paves the way for developing more trustworthy and effective systems for survey generation.
+9 Limitations
+We acknowledge several limitations in the current scope of this work.
+First, SurGE is limited to the Computer Science domain.
+This choice allows us to utilize the extensive open-access literature available on arXiv.
+More importantly, it aligns with our team’s domain expertise.
+This background enables us to rigorously validate the ground-truth surveys and ensure a standard of reliability that is difficult to maintain in other fields.
+Second, concerning evaluation methodology, assessing the quality of scientific literature inherently involves subjectivity.
+While our automated metrics provide a scalable and reproducible standard, they serve as quantitative proxies rather than perfect replacements for the nuanced judgment of human experts.
+Third, and perhaps most importantly, high-quality survey papers often go beyond summarizing existing work.
+Instead, they propose novel insights and identify future research directions.
+While such forward-looking analysis is a hallmark of the most influential surveys, SurGE does not currently evaluate this capability due to the significant challenges in establishing a standardized ground truth for “insightfulness”.
+We leave the development of metrics for these high-level cognitive tasks to future work.
+10 Ethical Considerations
+This work utilizes publicly available metadata from arXiv, strictly adhering to the Creative Commons CC0 1.0 Universal license and arXiv’s Terms of Use.
+We emphasize that the SurGE dataset consists solely of metadata and derived features; we do not host or redistribute original full-text PDF files, ensuring full copyright compliance.
+We also acknowledge the risks inherent in using Large Language Models for scientific writing, such as the potential for hallucinations and citation bias (e.g., over-representing specific authors or institutions).
+SurGE is specifically designed as an evaluation benchmark to quantify and mitigate these issues.
+By providing a rigorous testing ground for factual consistency, we aim to foster the development of trustworthy and responsible scientific writing assistants.
 Appendix A Reference Extraction and Processing
@@
 By using this set as the ground truth, we can quantitatively measure the coverage and recall of a system-generated survey’s references, providing a clear metric for its performance.
-Appendix B Details of the Academic Corpus
-Table 3 provides an overview of the metadata schema used in our academic corpus. Each entry is structured to support efficient retrieval and interpretability.
+Appendix B Data Field Descriptions
+In this section, we provide the detailed metadata schema for both the ground-truth survey dataset and the academic retrieval corpus utilized in SurGE.
+Table 5 details the pre-processed fields for each survey instance.
+These fields include not only standard metadata (e.g., Title, Authors) but also the parsed structural tree and the ground-truth list of cited documents, which are essential for evaluating hierarchical generation and citation recall.
+Additionally, Table 6 provides an overview of the metadata schema used in our academic retrieval corpus. Each entry is structured to support efficient retrieval and interpretability.
+Appendix C The Use of Large Language Models
+Large language models were used as a tool to help refine the language and improve the clarity of this manuscript.
+Following any such use, the authors reviewed and edited the content to ensure its accuracy and originality, and take full responsibility for the final text.
+Table 5: Description of metadata fields in the pre-processed survey dataset.
+Field Description
+SurveyID A unique identifier for the survey.
+Authors List of contributing researchers.
+Title The title of the survey paper.
+Year The publication year of the survey.
+Date The timestamp of publication.
+Category Subject classification following the arXiv taxonomy.
+Abstract The abstract of the survey paper.
+Structure Hierarchical representation of the survey.
+All_Cites List of document IDs cited in the survey.
 Table 6: Fields and Descriptions for the Literature Knowledge Base.
@@
 doc_id A unique identifier assigned for reference and retrieval.
-Appendix C The Use of Large Language Models
-Large language models were used as a tool to help refine the language and improve the clarity of this manuscript.
-Following any such use, the authors reviewed and edited the content to ensure its accuracy and originality, and take full responsibility for the final text.
-Appendix D Baseline Implementation
-D.1 Retrieval-Augmented Generation
+Appendix D Paper Retriever Training Details
+In this section, we provide the detailed formulation for the dense retriever described in Section 5.1.
+As previously mentioned, we leverage a dual-encoder architecture based on roberta-base.
+The training process utilizes the benchmark dataset, where each topic description $t$ serves as the query $q$.
+For a query $q$ and a paper abstract $d$, we construct their input representations by prepending the special token $\mathrm{CLS}]$ and appending $\mathrm{SEP}]$.
+Formally, let $X(q)=[\mathrm{CLS}]\,q\,[\mathrm{SEP}]$ and $X(d)=[\mathrm{CLS}]\,d\,[\mathrm{SEP}]$.
+We feed these tokens into the encoder to obtain the contextualized embedding of the $\mathrm{CLS}]$ token:
+$\mathrm{Emb}(X)=\mathrm{transformer}_{[\mathrm{CLS}]}(X).$ (7)
+The similarity score between $q$ and $d$ is computed as the dot product of their embeddings:
+$S(q,d)=\mathrm{Emb}(X(q))^{\top}\cdot\mathrm{Emb}(X(d)).$ (8)
+During training, each query $Q$ is paired with relevant documents $d^{+}$ from the ground truth to form positive pairs, while negative samples $d^{-}\in N$ are randomly sampled from the corpus.
+The retriever is optimized via the softmax cross-entropy loss:
+$\mathcal{L}(Q,d^{+},N)=-\log\frac{\exp\bigl(S(Q,d^{+})\bigr)}{\exp\bigl(S(Q,d^{+})\bigr)+\sum_{d^{-}\in N}\exp\bigl(S(Q,d^{-})\bigr)}.$ (9)
+This objective encourages the model to assign higher scores to relevant papers while minimizing scores for irrelevant ones.
+After training, the retriever is used to fetch the top-ranked papers for the subsequent generation stage.
+Appendix E Baseline Implementation
+E.1 Retrieval-Augmented Generation
 In the first baseline, we combine retrieval with a direct generation approach.
@@
 Although this baseline follows a straightforward two-step approach, it provides a clear assessment of how effective retrieval-based summarization can be when coupled with an LLM’s generative capabilities.
-D.2 AutoSurvey
-AutoSurvey (Wang et al., 2024d) implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions.
+E.2 AutoSurvey
+AutoSurvey (Wang et al., 2024) implements a multi-stage survey generation pipeline that starts with a high-level outline and proceeds through iterative expansions.
 We adapt it to use our fine-tuned retriever in place of its original retrieval mechanism and keep the number of retrieved references consistent for fairness.
@@
 Finally, all refined sections are merged into a coherent final survey, with transitions and citation references carefully aligned.
-The workflow iterates over these stages, leading to incremental improvements in both thematic coverage and presentation quality.
-D.3 StepSurvey
-StepSurvey (Lai et al., 2024b) is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases in a sequential manner.
+This process can be repeated for multiple trials, with evaluation feedback guiding further refinement and the best candidate selected.
+E.3 StepSurvey
+StepSurvey (Lai et al., 2024) is a more granular generation strategy that also begins with retrieving the top 100 candidate papers for a given topic $t$ but proceeds through distinct planning and drafting phases sequentially.
 This baseline is proposed by a team named “ID” in the NLPCC2024 competition task 6 (Tian et al., 2024b).
@@
 Subsequently, it uses the primary headings to guide the selection of secondary or finer-grained topics, each mapped to a relevant subset of the retrieved papers.
-The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging greater control and consistency across sections.
+The LLM then produces a full draft by writing each subsection with explicit attention to references and academic conventions, thereby encouraging consistency across sections.
 Throughout this process, the system attempts to maintain a balanced level of detail, striving for a clear exposition of important subtopics while avoiding excessive verbosity or redundancy.
 By structuring the content in incremental steps, StepSurvey aims to achieve coherent organization and thorough coverage of the literature.
-Appendix E Prompt Template
+E.4 SurveyForge
+SurveyForge (Yan et al., 2025) introduces a two-stage framework that combines heuristic outline generation with memory-driven content writing.
+To ensure a fair comparison, we strictly follow the official implementation [^3] https://github.com/InternScience/SurveyForge but replace the original retrieval component with the retriever described in Section 5.1, keeping the candidate document pool the same as other baselines.
+The pipeline begins by constructing a hierarchical outline, leveraging structural patterns derived from human-authored surveys to organize the topics.
+Subsequently, the method employs a Scholar Navigation Agent (SANA) to synthesize the content.
+In our adapted workflow, this agent queries the pre-retrieved papers to draft specific subsections, ensuring that the generated text is structurally coherent and grounded in the provided references.
+Appendix F Extended Related Work
+F.1 Retrieval-Augmented Generation
+Retrieval-Augmented Generation (RAG) has emerged as a cornerstone for scientific writing (Lewis et al., 2020; Gao et al., 2023; Su et al., 2024).
+Fundamentally, RAG pipelines follow a retrieve-then-read paradigm, identifying relevant documents to condition the generator (Guu et al., 2020).
+While this paradigm shares the same core logic as Survey Generation (incorporating external knowledge to generate content), standard RAG approaches struggle to meet the unique demands of survey writing.
+Specifically, naive RAG pipelines face three critical limitations in this context.
+First, standard RAG pipelines lack the hierarchical planning required to maintain a logical narrative flow, often treating the task as simple QA rather than document composition.
+Second, they fail to capture complex inter-document relationships (e.g., chronological evolution or methodological contrasts), which are essential for synthesizing conflicting information from massive corpora.
+Third, the limited context window in standard models often leads to information loss when processing hundreds of references.
+To address the structural complexity of full-length surveys, the field is shifting towards specialized Agentic Workflows that decompose the task into autonomous phases of planning, execution, and verification.
+We introduce this emerging paradigm in the following subsection.
+F.2 Agentic Survey Generation
+Before the advent of Large Language Models, scientific survey generation was often viewed as an advanced form of Multi-Document Summarization (MDS), yet it presents unique challenges regarding scale and structure Lu et al.
+(2020); Ma et al.
+(2022).
+Unlike traditional MDS, which focuses on distilling information into concise overviews (Nayeem and Rafiei, 2024; Bleiweiss, 2023), survey generation aims to synthesize extensive literature into a coherent, lengthy manuscript with a hierarchical structure Wang et al.
+(2024).
+Early approaches struggled with these constraints due to the limited context windows of language models.
+The advent of LLMs has triggered a paradigm shift from simple multi-document summarization to complex agentic workflows.
+For instance, AutoSurvey (Wang et al., 2024) pioneers a multi-stage pipeline that moves beyond simple drafting.
+It generates a structured outline and performs iterative expansions, explicitly refining sections to address factual and stylistic inconsistencies.
+Similarly, StepSurvey (Lai et al., 2024) employs a granular strategy, prioritizing sequential planning where secondary topics are mapped to specific paper subsets before drafting to ensure balanced detail.
+From another angle, SurveyForge (Yan et al., 2025) introduces a distinct approach by combining heuristic outline generation with a memory-driven Scholar Navigation Agent (SANA) to synthesize content.
+Most recently, SurveyX (Liang et al., 2025) introduces AttributeTree templates to distill reference information, and uses extraction and generation templates to automatically construct auxiliary figures and tables, followed by a re-polishing (post-refinement) process.
+Despite these architectural innovations, evaluation remains a critical bottleneck.
+Existing works predominantly operate within a "self-validation" paradigm, where metrics are either unscalable or tailored to specific pipelines.
+For instance, approaches like StepSurvey rely heavily on human evaluation (Tian et al., 2024a), which strictly limits scalability and reproducibility.
+Conversely, framework-oriented works such as AutoSurvey and SurveyForge propose custom metrics designed to validate their own specific workflows.
+This fragmentation creates a "metric bias" that obscures fair cross-system comparison, highlighting the urgent need for a standardized, method-agnostic benchmark to assess these diverse agentic strategies.
+F.3 Long-Context Evaluation and Long-Form Generation
+[…]
+In parallel, long-form generation has been widely studied under long-document summarization and long-form QA.
+Datasets such as GovReport Huang et al.
+(2021) and QMSum Zhong et al.
+(2021) require generating summaries from lengthy, information-dense sources, and BookSum Kryściński et al.
+(2022) focuses on narrative summarization with rich discourse and long-range dependencies.
+Despite their value, these tasks typically target a summary or answer and rarely require explicit bibliographic grounding, faithful claim-to-citation mapping, or field-level structural synthesis.
+Thus, scientific survey generation raises a different evaluation challenge.
+A survey is not only long, but it must also be academically valid, requiring (i) broad bibliographic coverage of seminal work, (ii) faithful attribution of claims to references, and (iii) coherent structural taxonomy that organizes and contrasts prior studies.
+This gap motivates a dedicated, method-agnostic benchmark and diagnostic evaluation framework that can fairly compare diverse survey generation pipelines.
+Appendix G Prompt Template
 This section provides the detailed prompt templates used for our LLM-as-a-Judge evaluation, as described in the main body of the paper.
@@
 Paragraph: {Paragraph}
-Appendix F Selected Ground Truth Survey
+Appendix H Selected Ground Truth Survey
 The following three tables provide the complete list of the ground-truth surveys in the SurGE benchmark.
@@
 Transformers in Vision: A Survey 2021 cs.CV 3276
+A Survey of Knowledge-based Sequential Decision Making under Uncertainty 2020 cs.AI 26
+Automated Machine Learning on Graphs: A Survey 2021 cs.LG 97
+Active Divergence with Generative Deep Learning - A Survey and Taxonomy 2021 cs.LG 23
+[…]
+A Survey on Multi-modal Summarization 2021 cs.CL 73
+Generative adversarial networks in time series: A survey and taxonomy 2021 cs.LG 105
+Towards Reasoning in Large Language Models: A Survey 2022 cs.CL 841
+Survey on reinforcement learning for language processing 2021 cs.CL 180
+Creativity and Machine Learning: A Survey 2021 cs.LG 56
 Image Classification with Deep Learning in the Presence of Noisy Labels: A Survey 2019 cs.LG 433
@@
 Opportunities and Challenges in Explainable Artificial Intelligence (XAI): A Survey 2020 cs.CV 1062
+Transformers in Medical Imaging: A Survey 2022 eess.IV 889
+A Survey of Privacy Attacks in Machine Learning 2020 cs.CR 368
+A Survey on Deep Reinforcement Learning for Data Processing and Analytics 2021 cs.LG 44
+A Survey on Dialogue Summarization: Recent Advances and New Frontiers 2021 cs.CL 111
+A Survey of Machine Learning for Computer Architecture and Systems 2021 cs.LG 91
+Inspect, Understand, Overcome: A Survey of Practical Methods for AI Safety 2021 cs.LG 78
+Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394
+A Survey on Recent Approaches for Natural Language Processing in Low-Resource Scenarios 2020 cs.CL 390
 Domain Adaptation and Multi-Domain Adaptation for Neural Machine Translation: A Survey 2021 cs.CL 108
@@
 Stabilizing Generative Adversarial Networks: A Survey 2019 cs.LG 149
-A Survey of Knowledge-based Sequential Decision Making under Uncertainty 2020 cs.AI 26
-Automated Machine Learning on Graphs: A Survey 2021 cs.LG 97
-Active Divergence with Generative Deep Learning - A Survey and Taxonomy 2021 cs.LG 23
-[…]
-A Survey on Multi-modal Summarization 2021 cs.CL 73
-Generative adversarial networks in time series: A survey and taxonomy 2021 cs.LG 105
-Towards Reasoning in Large Language Models: A Survey 2022 cs.CL 841
-Survey on reinforcement learning for language processing 2021 cs.CL 180
-Creativity and Machine Learning: A Survey 2021 cs.LG 56
 A Survey of Deep Reinforcement Learning Algorithms for Motion Planning and Control of Autonomous Vehicles 2021 cs.RO 92
@@
 Retrieval-Augmented Generation for Large Language Models: A Survey 2023 cs.CL 1951
-Transformers in Medical Imaging: A Survey 2022 eess.IV 889
-A Survey of Privacy Attacks in Machine Learning 2020 cs.CR 368
-A Survey on Deep Reinforcement Learning for Data Processing and Analytics 2021 cs.LG 44
-A Survey on Dialogue Summarization: Recent Advances and New Frontiers 2021 cs.CL 111
-A Survey of Machine Learning for Computer Architecture and Systems 2021 cs.LG 91
-Inspect, Understand, Overcome: A Survey of Practical Methods for AI Safety 2021 cs.LG 78
-Survey on Evaluation Methods for Dialogue Systems 2019 cs.CL 394
-A Survey on Recent Approaches for Natural Language Processing in Low-Resource Scenarios 2020 cs.CL 390
 DL-Traff: Survey and Benchmark of Deep Learning Models for Urban Traffic Prediction 2021 cs.LG 180